A Complete Bibliography of Publications in the SIAM Journal on Algebraic and Discrete Methods

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

19 October 2023 Version 2.09

Title word cross-reference

(0, 1, −1) [MW81]. (A, B) [Bha83]. (A, K) [CDS84]. (A ∘ A−T) [JS86]. 0 [BB87]. 1 [Ber83, BB87]. 12 [GM87a]. 2 [DH87, Tay82]. 22 [DF83b]. 2 × 2 [Uhl85]. 3 [AP86, HS85, Mon86]. 600 [LT80]. A [EJ82]. AAT = vI − J [MW81]. AX + X^{∗}A^{∗} = C [LR83]. Ax = b [EJ82]. b [Coo83]. P_{n}^{g}(A) [VC82]. D [CDJ82]. sl(2, C) [Pro82a]. GF(3) [MS81]. H [Har86]. J′ [Tod84b]. K [CDS84, ACP87, CN84, CK87, Hwa86, LL87, Mac81]. L [MW87]. l^{1} [CK86]. l_{1} [Rot84]. l_{∞} [Rot84]. l_{p} [DW80]. λ [JS81]. log [CT80]. LU [AO86, Dav86, FP86, Har86, Mat84, VC82]. M [AO86, Bar86, Buo86, FP84, Gam87, JOv83, Lew80, Ros84, Smi81, VC82, Gam86].

N [Dia80, Tan84, CW87, Hwa86, Kir83, Sal84]. N × N [MD86]. O((n [CT80]. O(n^{2}) [OBB81]. O(|V |^{2}) [HS85]. P [GJR83, GJ86, CT80]. p)^{2}) [CT80]. q [Dun80, Kir83]. QR [GM86, Chu84]. R^{n} [Tod84b]. S^{n−1} [Chu86]. t [GLL80, Moo83]. v × v [MW81]. Z_{2}^{k} [Mor86]. Z_{2}^{n} [Tan83].

-ary [Kir83, Moo83]. -cell [LT80]. -center [CT80]. -chain [Ber83]. -circulant [MW81]. -connectivity [Tay82]. -cube [Sal84]. -cut [HS85]. -decompositions [Har86, Mat84]. -designs [GLL80]. -distance [DW80]. -domination [CN84]. -edge-connected [CW87]. -functions [MW87]. -gons [Tan84]. -Hahn [Dun80]. -invariant [Bha83]. -Jacobi [Dun80]. -matching [Coo83].

-matrices [AO86, Buo86, Har86, JOv83, Ros84, Smi81, VC82]. -matrix [Bar86, FP84]. -monotone [JS81]. -NN [Mac81]. -out-of- [Hwa86]. -permutation [LL87]. -person [Dia80]. -recursiveness [GJR83, GJ86]. -sequences [Gam86, Gam87]. -stability [CN84]. -stable [CDJ82]. -tree [ACP87]. -trees [AP86, CK87].

2* [APSZ81].

abnormal [KS87b]. abstraction [Cha83]. Acknowledgment [MMP80a]. Acyclic [Gan81, BE84, BS86]. addition [AMT86]. additive [GS80]. Adjacency [Gil81, Wal86]. adjoint [HU85]. adjugate [HU85]. Aggregation [ANW84]. aided [Gol85a]. Algebraic [SW87, Gam87, GP83, SCC83, WW83, WW87, Wim82]. algorithm [Bar82, BS87, BL86, Car84, CBS85, CT80, Chu84, DS81, DS87, HS85, KV86, OBB81, Phi87, de 87]. algorithmic [Abe85, LPHH84, LN80]. Algorithms [ES82, Bac82, BCCL83, EHH87, HT86, Min85b, Sax80, Szp87, Vic84, Whi86, WS82, CP87]. alignments [GHW86]. allocation [ANW84, Lan84]. almost [HS86]. alphabetic [KS81]. alternating [GH87a]. among [Win83]. Amortized [Tar85]. analog [CH87]. analogue [Far86]. analysis [GLM81, Hav86, Ple86, Rot81, Shr84, Szp87, Won80, Zem85]. antichains [Dem80]. antirectangles [Bou84]. aperiodic [dK82]. application [BP85, CBS85, GJR83, Liu87, MIN87, Sam80, SCC83, Tog86]. applications [Bar86, Bar84, CLR87, CNW80, Dre82, EHH87, KKP85, MN80]. applied [CBS85, Ple86, SW87]. approach [Abe85, Ber80a, CK87, CT82, FS84, GP83, RP85, Rot81]. approximate [EJ82, Min85b]. Approximation [BCCL83, HR86, Hav86, Her84a]. approximations [Gup86]. arbitrary [Her84a]. arc [OBB81]. arcs [Bos84, GJMP80]. arising [BJW85, GGH80, Mat84, Szp87, Tam83]. arithmetic [AW80]. arrangement [Len82]. array [JS86, MD86]. art [Gyö86]. ary [Kir83, Moo83]. aspects [AB84, Her84b, JS86]. assay [BH85b]. assignment [RP85, WCLO86]. assignments [DH87]. associated [Sak82]. association [Sam80]. Asymptotic [BP85, Dia80, BC81, Bul85, FW80, RP85, DGK80]. automata [Tch85]. automation [PR82]. average [Rus80]. averages [Mor86].

balanced [BL86, HKS85, Tam83]. ball [Rot84]. Balloting [WCLO86]. band [EGL87, Per87, Tre85, Tre86]. bandwidth [APSZ81, Chu85b, MPS85, Mon86, Sax80]. Banzhaf [Bol83]. bar [WW87]. bar-and-body [WW87]. based [LLM84, SS85]. bases [GS80]. basis [Bar87, GH87b, Pro85]. be [RV83]. behavior [Bul85, DGK80, Gol82]. behaviour [Gol85b]. belonging [HHS82]. benefits [PR82]. best [EJ82, Min85b]. between [CK86, GH82]. bi [Wim84]. bi-infinite [Wim84]. Bicycles [Ber86]. bin [BC81, Gal86, She81]. bin-packing [BC81, Gal86]. binary [CW82, CH80b, KS81, Meh86, Rus80, Tch85]. binomial [Meh86]. bins [CGJ82]. bipartite [HK86]. Bisection [GW85]. Bithreshold [HM85]. bivariate [CS87]. block [CCPR82, Mat84, MIN87, Ple86, von81]. block-triangularization [MIN87]. blocking [Cha87, LN80]. body [WW87]. books [CLR87, Hea87]. Boolean [Buz82, Hul84]. Bott [Cha83]. bound [BC81, Gal86, HR86, Pip80]. boundary [HHW81]. bounded [CG82, Fis83]. Bounds [BP82, EOZ87, PP85, BH85a, Kör86, Kri85, Len82, Mey80, KS87a]. branching [BT81]. Broadcasting [FP81]. BVP [Mat84]. Byers [CT15]. can [HS86, RV83]. Candidate [Dem80]. Canonical [CP83, CNW80, MIN87]. capacity [Far86]. capital [Chu85a, Jon82]. cardinality [VC82]. Carlo [Gra81]. Cascade [AMT86]. case [CK86, Gag83, GH82, Hil87, Uhl85, Won80]. cases [Per83]. caterpillars [APSZ81, Mon86]. cell [LT80]. center [CT80, Chu84]. central [Gup86]. centroid [Mer86]. certain [CS84, FP84]. chain [Ber83, HR86, Hav86, Mey80]. Chains [GM86, BP86, Bar86, Hey87, KMS84, Ste86]. Change [Bar87]. channel [CGW87, MMP80a, MMP80b, RP85]. characteristic [BFHMV84, Ran85]. Characterization [AP86, Bar80, HM84, AW80, CDJ82, SCC83]. Characters [MP84b, KV86]. chemical [Shr84]. Chevalley [Sta80b]. choice [Fis84]. Cholesky [Sam80]. Chopped [Per83]. chord [SCC83]. chordal [CN84, GRE84]. chords [GJMP80]. circle [GW85]. circuits [AT85]. circulant [BW86, MW81, Wat87]. circular [GJMP80, OBB81]. circulations [GP85]. class [Bha82, Bol87, Dia80, JOv83, RZ82, Sak82, She82, Tam83, Tre86]. classes [DGK83, KCP82]. classical [Bar84, Had85]. claws [Taz85]. climbing [DS87, Tov85]. closed [RV83]. clustering [Her84a, Her84b]. coalitional [LM82]. code [CDS83]. codes [CW82, Gag81, Gag83, KS87a, KS87b, MS81, Mol86, Phe83, Phe84, Tie85, von81, KS87b]. Coding [CTPR85]. coefficients [Bul85]. color [Bou84]. Coloring [CCPR82, dPR82, AG84, CC86, GJMP80, HK83, OBB81, PP85, PS86]. colorings [Gil80a, GW85]. colouring [KKP85]. Column [Dav86]. Combinatorial [AB84, MIN87, Cha82, DH82, GGH80, Phe83, TZ83]. combinatorics [Lin86]. combined [FP84]. commodity [ST80]. communications [LL87]. commuting [Per87]. comparison [BJW85, KS85]. comparisons [Ber80a, Hey87]. compartmental [Wal85]. competition [Ops82]. Complementarity [Tod84a]. complementary [BS86, Chr83]. complements [Pro87]. complete [CH80b, HK86, Taz85, Wim82, CCPR82, GJ83, Mon86, Yan81]. completely [HR86, Hav86, KMS84]. completeness [DK87, Gil80b]. complex [CH87, Had85, LLM84, Uhl85]. Complexity [ACP87, Wal85, CM84, GJMP80, Gat83, LPHH84, Len82, MS86, Tar85, Yan82, CP86]. comprehensive [HR85]. compressible [BHP86]. computation [Ebe87, KKS87, Ops82]. computational [Tar85]. Computations [KOY86, Ple86]. Compute [GM86, PR85]. computer [BP85, Gol85a, Lan84]. Computing [BFHMV84, DG86, GH87b, Jon82, Yan81, BP86, FP86, Hey87, KMS84]. concatenations [KSSS80]. concentrators [Tan84]. concept [SCC83]. condition [Mey80]. conditional [Gra81]. cones [BLP87, – Dok87]. congested [AM81]. conjecture [She81]. connected [CW87, RZ82, Tch85]. connection [CH80b, KKP85]. connectivity [BW86, FJ81, IW81, Tay82]. consecutive [DH87, Hwa86, SW86]. consecutive- [DH87, Hwa86]. consequences [Bha82]. consimilarity [HH86]. constant [JR85]. constrained [Cha83, Pro87, Tay82]. constraint [BL86]. construction [DS87, Gam87, KW80a, Mol86, Phe83, Phe84]. constructions [von81]. continuous [CT80, FW82]. contraharmonic [AMMT87, GM87b]. control [Bha83, Bol85a, PR82]. controllability [Bha83, CDS84, Wim82]. Controlling [AS80]. convection [Gup86]. convection-diffusion [Gup86]. Convergence [Joh85, Neu86]. Convergent [BP86, Ros84]. Convex [BH85b, JR85, BM84, GH82, KT81]. Convexity [FJ86]. convolutional [von81]. convolutions [EH87]. correcting [Phe84]. Correlation [Win83]. correspondence [Vo81]. cost [GH82]. costliest [KS81]. count [LAP82]. counterexample [She81]. Counting [Bra82]. coverage [MZH83]. Covering [AO81, AT85, CKSS81, CN85, BL86, FW82, Hoc83, KS87a, KS87b]. coverings [EOZ87, WT81]. criteria [Her84a]. critical [EZR85]. critically [CW87]. crosscorrelations [Gam86]. Crossing [GJ83]. cube [Mat87, PP85, Sal84]. current [Ber83]. Curves [Chu86]. cut [HS85, Len82, Sal84]. cutwidth [Chu85b]. cycles [BW80, GS84, dK82]. cyclic [CC86].

DAC [PR82]. data [BP85, CRS83]. decision [BS87, DK87]. decoding [BS87]. decomposable [HR86, Hav86, KMS84]. Decomposing [BW80]. Decomposition [Cun82, Has85, Taz85, AG84, CS87, CH87, Chu81, DER85, Sam80, Tog86]. decompositions [Bou87, Har86, Mat84]. Decreasing [BC81]. defectives [Chr83]. defining [Coo83]. definite [HM84]. degrees [GJ86]. demand [FW80]. density [Ber87]. dependent [FW80, NS81]. depth [PY82]. description [Uhl85]. design [CLR87, Gol85a, Sug83, Won80]. designs [CMW85, CCPR82, GLL80]. detection [Chr83]. Determinacy [Pro83]. determinant [Ber80b]. determinants [Bul85]. Diagonal [BPP81, BH83, CDS83, ES82, HMS87, HH86]. diagram [LLM84]. dictionary [SW86]. Difference [Cla86, Gup86]. different [HS81]. differential [Ber87, BJW85, CP83, Cam83, Cla86, Grü86]. Differentiate [GM86]. diffusers [Buc86]. diffusion [GGH80, Gup86]. digital [AS80, Joh85]. digits [Pro82b]. Digraph [Bou87, KS85, SP84]. digraphs [Gan81]. Dilworth [Sak80]. dimension [CL84, EOZ87, KS87b, Wim84, Yan82]. dimensional [ABBW87, CGJ82, HG85]. dimensions [BS83]. direct [FP84, Hey87, WT81]. direct-iterative [FP84]. directed [BE84, BH85c, Cun82, Gil80b]. directional [FJ81]. Discounted [ACN80]. Discrete [Fis84, Per87, Bal85, Bos84, CS87, Cia85, Dre82, FS85, Grü81, Per83, TS85, Vic84]. discretization [Grü86]. Disjoint [RS85, CH81, EOZ87]. Disjunctive [Bal85]. displacement [BM87]. distance [DW80]. distances [MT83]. distinct [KW80a]. distributed [BM80]. Distribution [CGW87, HR86, Hav86, PR82, SW86]. distributions [Bar80, BP86, CS87, FP86, Hey87, KMS84]. domains [AJ84]. dominant [Neu86]. dominating [CK87]. domination [CN84, LPHH84]. Double [CR82, Pro85]. double-basis [Pro85]. Doubly [HG85]. Doubly-periodic [HG85]. Drazin [Bra82]. dual [BHP86, Coo83, Fis87, Hal85]. duality [HM84]. Duffin [Cha83]. duplicate [Wes82]. Dynamic [Sax80, BL86, CK87, HR85, Rot81, Yao82]. Dynamic-programming [Sax80]. Dynamical [Gol85b]. dynamics [GH85].

edge [CW87, PS86]. edge-coloring [PS86]. effect [dN85]. effective [Buz82]. effects [AG83]. Efficient [HT86, LN80, PS83, PS86, Bol87, Lam87]. Eggenberger [BP85]. eigenvalue [AG84, SN85, Tre86]. eigenvalues [Chu86]. Eigenvectors [Grü81, dN85]. elections [Bol83]. electricity [Ber83]. elements [MP84b]. elimination [Gil80b, MN80]. Embedding [CLR87, Hea87]. Embeddings [ABBW87, ACP87]. encoding [MN84]. Ensembles [KW80b]. entries [BJOv87, SW86]. Entropy [FB84, EGL87]. enumeration [BW80, Lin86, LC84]. enumerators [Gag81, Gag83, MS81]. enveloping [Gup86]. equality [Hil87]. equation [Ber87, BJW85, FS85, Grü86, LR83, MW81, Szp87, TS85, Wim82]. equations [BJW85, Buz82, CP83, Cam83, Cla86, Gat83, GGH80, Gup86, MIN87, Ran85, RZ82, UP83, Vic85]. Equilibria [AM81, Dia80]. Equilibrium [Zem81]. Equivalence [DGK83]. equivalents [de 87]. Erdos [DF83b, FF80, Sta80b]. Erdos-Ko-Rado [DF83b, FF80]. ergodic [Bar86, FB84]. Erratum [Red86]. Error [Kri85, AG83, HR86, Hav86, Phe84]. estimate [GM86]. estimates [Mac81]. Estimation [NR83, CC86]. Eulerian [Bou87, GP85]. evaluated [RV83]. evaluating [PR82]. events [BM80]. Exact [UP83, Vic84]. Examples [EZR85]. excavations [DM86]. exchange [LLM84]. existence [GM87a]. Expanders [Buc86]. expansions [Per83]. Expected [KT81]. Explicit [Rot84, Tan84, Tre85, Uhl85]. exponent [Ros82, Sha87]. exponential [LN80]. exposed [BLP87]. extension [Ber87, Min85a]. extensions [CFG80, DD85, EGL87]. extractions [Gat83]. Extremal [Cha87, GW80]. Exxon [Dre82].

facets [AJ85, Buc87]. facilities [HS81]. facility [DW80]. Factorization [GM86, AO86, Buo86, Dav86, Lam87, Liu87, Tch85, VC82]. factorizations [DS87, FP86, Ste86]. factors [Rub86]. failures [BH86]. Families [FK84, CRS83]. Farey [Col82]. Fast [KKS87, DS81]. faulty [DC81, HRS87]. fewer [KKK83]. field [Uhl85]. fields [BFHMV84, Gol80, SL83]. fill [Wen87, Yan81]. fill-in [Wen87, Yan81]. Finding [MT83, ACP87, DS81]. finish [DFL82]. Finite [LM82, ABBW87, BFHMV84, DD85, FFF^{+}86, Gam86, Gol82, Gol80, Lih82, Mey80, MP84b, SL83]. finite-dimensional [ABBW87]. Fit [BC81]. Fixed [Gol82, FS84]. fixpoints [SS85]. FKG [She80]. flights [Tak81]. flow [BHP86, Chu84]. flows [de 84a]. Fluctuation [NS81]. fluid [BHP86, GH85]. forests [DW85, GJTY83]. form [Bol87, HH86, MIN87, RV83]. formal [LC84]. forms [CP83, KKS87]. formula [Ber80b, Buz82, LR81]. formulae [Gra81]. formulas [Ber83, Tre85]. formulation [Cla84]. formulations [GM87b]. Four [GM87b, JM87]. Fourier [Per81]. fractional [Hoc83]. frameworks [WW83, WW87]. Fredman [Kör86]. Fredman-Komlós [Kör86]. free [CN84, GM87a]. full [BL83]. function [Bol87, Bul85, DH82, GJR83, Mol86, Pro82b]. functionals [DGK81b]. Functions [FK84, DGK81a, FW80, Gol82, Grü81, Her84a, Her84b, KL84, MW87, PS83]. Further [Hey87]. fuzzy [Has85].

gain [JS86]. galleries [KKK83]. gallery [Gyö86]. game [Hwa84, Rab85]. games [ACN80, Bol87, Dia80, GH82, LM82, Owe86, SS85]. general [Phe84, Pro85, Wen87]. Generalized [Bha83, DGK81a, Meh86, BP85, Bar84, BR82, CN85, Chu84, HSD81, KKP85, Liu87, Mol86, Tan84, Tie85]. Generalizing [Pro82b]. generated [Bul85]. Generating [Wal83]. generic [LY82]. geometric [Gol85a, Zem85]. geometry [Gam86, Lin86, WW83, WW87]. given [Buc87]. gons [Tan84]. Gossiping [Wes82, HRS87]. graceful [BH85c]. Graph [AG84, GLM81, GH85, BE84, Bar82, Ber80a, Far86, GW80, HHS82, HVW82, JM87, KKP85, LAP82, LLM84, LC84, Mer86, Ops82, Owe86, Taz85, de 84b]. graph-colouring [KKP85]. graph-restricted [Owe86]. graph-theoretical [Ber80a]. graphs [BH85c, BW86, CN84, Chu81, CLR87, Col82, CR82, CL84, CW87, Cun82, Dji82, DH87, FJ86, Fin85, FJ81, GS84, Gil80b, GRE84, GJ86, GS80, GLM83, GLM84, HM85, HT86, Hea87, HK86, HK83, JOv83, KSSS80, MW83, MN84, OBB81, PS86, Sak82, SK86, Sax80, SS85, She82, SCC83, Tay82, Wan83, Wan84]. gravity [Elf81]. Greedoids [KL84]. greedy [BL86, FW82, HKS85]. Group [EH87, GM86, AJ85, Bac82, CH80a, CH81, DK87, Fis84, HHW81, HSD81, Hwa84, KV86, Lam87, Meh86]. groups [MP84b, Sta80a, Sta80b]. grown [MP84a]. growth [FS80, Rot81].

Hadamard [BS87]. Hahn [Dun80]. hair [Mon86]. hairs [APSZ81]. Half [DGK81b]. Half-plane [DGK81b]. Hamiltonian [GS84]. Hammond [GJR83]. handicap [Ber80a]. Hard [Lin86, Bou84, Sta80a]. harmonious [GS80, HK83]. Hash [FK84]. having [GH87a]. Helly [Pec81]. Hermite [KKS87]. Hermitian [DGK83, Gag83, JR85, dN85]. Hessian [CC86]. heuristic [FW82, Lin87]. heuristics [Lan84, Won80]. hierarchical [Toc84]. hierarchy [Bal85]. Hill [Tov85, DS87]. hill-climbing [DS87]. homogeneous [FW80]. Homology [Sta85]. homometric [RS82]. hookup [KCP82]. Hopf [Hil87]. HSD [AMMT87]. Hypergeometric [HSD81]. hypergraphs [FJ86]. hyperplane [Rot84]. hypotheses [SSM84].

identifying [DM86]. II [CP87, DGK80, Gag83, Kan82a, KS87b, Red85b, Red86, Ros82]. ill [AG83]. ill-posed [AG83]. images [Joh85]. immanantal [Mer86]. Implementation [Pro85]. implications [BH83]. improve [Phi87]. Incidence [LR81, Sak80]. including [Tre86]. Incomplete [Buo86]. Incremental [CBS85]. indecomposability [BL83]. independence [BP82, Hul84]. independent [BE84, Wil86]. Index [Cam83, Bol83, Bra82, DG86]. indifference [CR82]. industrial [Tog86]. Inequalities [DF83a]. inequality [BR82, DER85, Hil87, She80]. inertia [JR85]. infinite [KV86, Wim84]. information [Kör86]. Inherited [BJOv87]. Instability [Vic84]. Integer [BT81, FF80, Sta85]. integral [Coo83]. intersecting [FFF^{+}86, Fis87]. intersections [Rot84]. interval [Fis81, Fis83, GW80, LAP82]. Intrinsic [Abe85]. invariant [Bha83, Chu85a, – Dok87, Jon82, SK86]. inventory [FW80]. Inverse [AT86, BJOv87, Cha83, DGK80, Lam87, Toc84]. inverses [MW87, Smi81, Tre86]. Inversion [GM86, Min85a, Tre85]. Invertible [EGL87]. Inverting [GLM84]. irreducibility [BL83]. irreducible [Gol80, Har86]. Isolating [AG83]. isomorphic [Taz85]. Isomorphism [KCP82]. iterations [BP86]. Iterative [KMS84, FP84, Gup86, Had85, Kri85, Ple86, UP83].

Jacobi [Dun80, Ebe87, Lew80]. Jacobian [NR83]. jointed [DM86]. jump [EZR85]. jump-critical [EZR85].

Keldysh [Ste86]. Kerdock

[Kan82b, Kan82a]. kernels [SS85]. keyblocks [DM86]. keys [Dem80]. Kirchhoff [Cha87]. KLYM [DF83a]. Knuth [KV86]. Ko [DF83b, FF80, Sta80b]. Komlós [Kör86].

Labelled [GJ86]. labelling [WCLO86]. large [AJ84, BW80]. largest [KW80b]. Latent [Fis87]. later [DF83b]. Latin [GM87a]. lattice [BS83, Gil80a]. Lattices [Neu83, CS84, Vo81]. Laurent [Bul85]. layered [MIN87]. layout [CLR87]. Layouts [LLM84]. lead [Chu86]. leading [BH85a]. Leaky [Ber83]. least [DGK80, MT83]. least-distances [MT83]. least-squares [DGK80]. Lefschetz [Sta80a]. Legendre [Grü86]. lemma [Min85a]. lemmas [MN80]. length [APSZ81, Cha87, Mon86]. level [DW85, Moo83]. lexicographic [Vo81]. Lie [MP84b]. limitations [Abe85]. limited [PY82]. limiting [Mey80, Per87]. line [BW86, Car84, DER85, Gal86, HS81]. line-connectivity [BW86]. line-sum-symmetric [DER85]. Linear [BL83, KOY86, Bol85a, Buz82, Cam83, CFG80, DD85, – Dok87, FP84, GP83, Had85, KL84, Len82, MIN87, OW85, Pro85, Pro83, Pro87, SSM84, Szp87, UP83, de 87, de 84a]. linear/nonlinear [MIN87]. lines [HRS87, MT83]. lives [Bha82]. Local [Mac81, Tov85]. location [CNW80, CT82, DW80, HS81, MZH83, Tam83, Zem85]. logarithms [BFHMV84]. lot [FW80]. lot-size [FW80]. low [KS87b]. lower [Gal86, Len82, Pip80]. Lyapunov [CT15, BN87, CDJ82, FS85].

majority [AJ84, Abe85]. Majorization [Lih82]. manifold [Chu84]. manipulability [Tog86]. Mappings [AJ85]. maps [BB87, DD85, FB84, Sak80, Wal83]. marketing [Zem81]. Markov [BP86, Bar86, BB87, FS84, FB84, GM86, HR86, Hav86, Hey87, KMS84, Lat82, Mey80, NS81]. Markov-dependent [NS81]. matched [GHW86]. matching [Coo83]. matchings [Vo81]. Mathematical [JS86, Dre82]. mathematics [Fis84]. Matrices [HMS87, AO86, AMT86, AT86, AMMT87, Bra82, BL86, BL83, BS86, Buo86, BB87, CDJ82, CC86, DER85, DGK83, EJ82, EGL87, ES82, Gan81, GH87a, GLM81, GLM83, HM84, Har86, Has85, HKS85, JS81, JOv83, JR85, KKS87, KW80a, Lew80, LR81, Mat84, MIN87, NR83, OW85, Per87, RZ82, Ros84, Sha87, SP84, Smi81, Tam83, Toc84, Tre85, Tre86, VC82, Wal86, Wat87, dN85]. Matrix [BH83, RV83, Bar86, BJOv87, BPP81, BJW85, CDS84, Cha82, Chu86, DGK81a, DGK81b, EH87, FS80, FP84, Grü81, HU85, HH86, LR83, Liu87, MN80, Min85a, Ros82, Ste86, TS85, Uhl85, Vic85]. matrix-valued [DGK81a, DGK81b]. matroid [Cha83]. matroids [BM84, IW81, Tod84a]. maximal [BE84, BR82, Col82, Wil86]. maximization [MS86]. maximize [DFL82]. maximizing [BCCL83]. Maximum [Fis81, HHS82, MZH83]. mean [AMMT87, Bha82, GM87b]. means [AW80, AT86, Chu86]. merging [HH82]. meromorphic [Bul85]. message [Dav86]. message-passing [Dav86]. metagraphs [Čul84]. metapaths [Čul84]. method [AO86, BHP86, DM86, FP84, Hal85, Kri85, Neu86, PR85, Pro85, SN85]. methods [Cla86, Ebe87, GLM81, Hey87, Hul84, KMS84, LC84, PR82, SW87, UP83]. metrics [CK86]. middle [Sal84]. middle-cut [Sal84]. min [Len82]. min-cut [Len82]. minimal [Coo83, PR85]. minimax [DW80]. Minimean [HS81]. minimization [BPR82, DGK81b, Mon86]. Minimizing [DH82, GP83]. minimum [CGW87, CW87, DFL82, GH82, Lin87, Wen87, Yan81]. minors [Cha82, GH87a]. mixed [MIN87]. Mixing [Mat87]. model [AB84, BP85, Elf81, Gol85a, HR85]. models [FW80, Pro87, SSM84, Wal85]. modes [LL87]. moments [SWG83]. monomial [Wat87]. monotone [JS81]. monotonic [PS83]. Monotonicity [Bol85b, GYY80, She80]. Monte [Gra81]. most [MP84a]. motions [WW87]. Muller [Tie85]. multi [Vo81]. multi-subset [Vo81]. multicandidate [Bol83]. multicommodity [ST80]. multigraphs [AT85]. multipartite [Taz85]. multiple [Cia85, Doi87, Tov85]. Multiplication [Bar84]. multiplicative [Rot81]. multiplier [DM86]. multipliers [Red85a, Red85b, Red86]. Multiplying [CW82]. multipole [FP81]. multiport [Cha87]. multiprocessor [Dav86, DFL82]. Multisets [dK82]. Multisplittings [OW85]. multivariate [Ber87]. Munkres’ [CBS85].

Nash [CT15]. Navier [Hal85]. near [Fin85]. near-traceable [Fin85]. nearest [FJ81]. nearest-neighbor [FJ81]. nearly [HR86, Hav86, KMS84, Phi87, Ros82, Sha87]. neighbor [FJ81]. Neighborhoods [Neu86]. Network [Col87, AM81, HS81, SW87, WS82, Won80, de 84a]. networks [Cha87, Gol85b, KKP85, MMP80a, MMP80b, Pip80, PY82, Pro83, ST80, Tch85]. neural [Gol85b]. Next [BC81]. Next-Fit-Decreasing [BC81]. nilpotent [Gan81]. NN [Mac81]. no [HHS82, Rab85]. nodes [Bar82]. non [SCC83]. non-series-parallel [SCC83]. nonabelian [AJ85]. nondestructive [BH85b]. nonisomorphic [Wal83]. nonlinear [Joh85, MIN87, Whi86]. Nonnegative [BJW85, EJ82, JS81, Sam80, BL83, DER85, FS80, Lew80, Vic85]. Nontestable [SSM84]. norm [MS86]. normal [Ber87, KS87b]. normality [BP85]. normalized [Gag81, Gag83, KS87a]. norms [Rot84]. Note [CT15, Chu85a, Gil80b, SP84]. nowhere [HVW82]. NP [CCPR82, GJ83, Gil80b, Mon86, Yan81]. NP-complete [CCPR82, GJ83, Mon86, Yan81]. NP-completeness [Gil80b]. null [CP86, CP87, GH87b]. number [BCCL83, BE84, Buc87, Buz82, Chu85a, Cia85, GJ83, GW80, KT81, Ops82, Pip80, Rub86, Wil86]. numberings [MD86]. numbers [Moo83, Sak80]. Numerical [Hal85, Cla86].

objective [KL84]. objects [Cia85]. Obtaining [Gol80]. occurrences [Kir83]. odd [Boo82]. on-line [Gal86]. one [DS87, Ebe87]. one-factorizations [DS87]. one-sided [Ebe87]. ones [BS86]. Operator [CT15, BN87, LC84]. operators [– Dok87, GM87b, Per87]. opposing [DW85, GJTY83]. Optical [Shr84]. optima [Tov85]. Optimal [CMW85, Chr83, DH87, Gra81, HSY85, MD86, BS83, BS87, CDS83, Jon82, SP84]. optimization [Bal85, BT81, CM84, Had85, HT86, Hul84, PS83, Rot84]. Optimum [SN85]. Order [DD85, Boo82, MMP80a, MMP80b, MP84b, Tie85, Yan82]. ordered [EZR85, GP83, Lih82]. orderings [KS81]. orders [CFG80, DW85, Fis81, Fis83, GYY80, Sak80, She80, WT81, Win83]. ordinary [BJW85]. orientations [GP85]. oriented [BM84, Tod84a]. originators [FP81]. orrery [AB84]. Orthogonal [Dun80, Bar84, Bar87, BPP81, CH87, MS81, Per83]. orthogonal-symmetric [CH87]. orthonormal [Phi87]. other [Bol85b]. outerplanar [Col82, Hea87, PS86]. overlap [SW86]. overlapping [Her84a]. overload [AS80]. overrelaxation [SN85].

packed [BCCL83]. packing [BC81, CN85, CGJ82, FW82, Gal86, HS86, She81]. Packings [HK86]. pair [KSSS80]. paired [Ber80a, KS85]. pairs [CDS84, CTPR85]. paradoxes [Bol85b]. Parallel [Whi86, WS82, Ebe87, HT86, KKS87, OW85, Ple86, SCC83, de 87]. parameter [Bha83, SN85]. Parametric [Gal86]. parametrization [DGK83]. parity [Mol86]. partial [AP86, CFG80, GYY80, Sak80, She80, WT81, Win83, Yan82]. partially [Lih82]. partition [Bol87, CM84]. partitioning [Bar82, Dji82, HSY85]. Partitions [Tan83]. passing [Dav86]. path [CG82]. paths [RS85]. patterns [CH80b]. pebbling [Liu87]. Perfect [CRS83, FK84, Mol86, Phe83, Sak82, She82]. perfection [CN85]. Performance [Lan84]. periodic [HG85, Zem81]. Permutation [Tch85, BW80, LL87, MP84a]. permutationally [GH82]. Permutations [BM87]. person [Dia80]. personnel [WCLO86]. perturbation [Bol85a, Mey80, dN85]. phase [Lat82]. phase-type [Lat82]. picycles [Fis83]. pivoting [Dav86]. placement [CBS85].

Plackett [Ber87]. Planar [DGK80, Dji82, HS85, Sak82]. plane [DGK81b, Gil80a, LLM84, LY82, Sug83]. planes [Kan82b, Kan82a]. plant [CNW80, CT82]. point [FS84, Gol82, HVW82, Lat82]. point-symmetric [HVW82]. polar [CH87]. Polya [BP85, Gol85a]. Polya-Eggenberger [BP85]. polyhedra [AJ81, Buc87, Cha87]. polyhedron [Coo83, Gil81, KT81]. polymatroid [BT81]. polynomial [BP82, Gat83, KKS87, Mer86, Per83, Ran85, SK86, Sax80]. polynomials [Bar84, Bar87, DGK80, Dun80, Gol80, PR85, Rub86, Ste86, Wan83]. polytopes [BM84, Tak81]. posed [AG83]. poset [DD85]. posets [DF83a, Pro82a]. positive [Bar86, CDS84, HM84]. postman [Gil81]. powers [FS80]. Practical [Dre82]. predator [RZ82]. Preemptive [de 84a]. prefix [SW86]. preserving [BL83, DD85]. presidential [Bol83]. prey [RZ82]. prey-predator [RZ82]. primal [CT82]. prime [Gat83]. primitive [Ros82, Sha87]. principal [BJOv87, GH87a]. priority [MMP80a]. Probabilistic [Zem85]. Probabilities [FFF^{+}86, GM86, Mey80]. probability [CS87, KV86]. probable [MP84a]. problem [Boo82, Bos84, Bum81, CG82, CT80, CH80a, CH81, CLR87, Cla84, CC86, CP86, CP87, CK87, CT82, Dji82, GS84, GGH80, Hoc83, HS85, HS86, HHW81, KS87a, KS87b, Lan84, Len82, MS86, MZH83, Mon86, Pro85, RP85, SN85, Tre86, Wen87, Won80, Yan82]. problems [ANW84, AT86, AJ81, AJ85, AG83, Bal85, BH85a, BT81, BH86, Bol85a, BL86, BHP86, CN84, CNW80, ES82, FW82, Hal85, JM87, Lin86, LC84, Rot84, SW87, Sug83, Tam83, Whi86, WS82, Zem85]. procedure [CBS85]. process [Lat82]. processing [CBS85]. processor [DFL82, Phi87]. product [Phe84, Sak80]. products [Bar87, RV83, WT81]. Profile [DW85]. programming [Bal85, BL86, CK87, Dre82, HM84, HR85, Pro85, Sax80, Sta85, Yao82, de 84a]. programs [FS84]. projection [CS87]. Projectionally [BLP87]. projective [Gam86]. prolate [Grü81]. proof [Ber80b, Cha82, Gyö86]. proper [OBB81]. properties [BW86, BL83, GYY80, KV86, Mac81, She80, de 84b]. property [Grü86, Pro82a, Sta80a]. proportional [Bol85b].

Quadratic [– Dok87, BJW85, CW82, DGK81b, HM84, Rub86]. quadrics [Gam86]. qualitative [GLM81]. quantizer [BS83]. quasidistance [SP84]. Quasimonotonic [EHH87]. queues [LN80].

R [CT15]. radius

[BS86, BB87, KS87a, KS87b]. Rado [DF83b, FF80, Sta80b]. Radon [Mor86]. Random [Gil80a, LT80, Tak81, BE84, Bar80, Buc87, FFF^{+}86, KT81, KW80b, MP84a, Mat87, MMP80a, MMP80b, Sam80]. Randomly [Fin85]. rank [LR81]. Ranking [KS85, Ber80a]. rankings [CK86]. rates [Mat87]. ratio [ACN80]. reaction [GGH80]. reaction-diffusion [GGH80]. real [Rub86]. realizations [SP84]. Rearrangeable [PY82, Pip80]. recognition [AP86]. recognizing [Sax80]. Reconstruction [KSSS80, Per81]. rectangle [BCCL83, FHM80]. rectangles [BH85a, CKSS81]. Rectangular [GLM83, GLM81]. Rectilinear [FHM80, Gyö86, Sak82]. recurrence [BPR82, Szp87]. recurrences [HG85, WS82]. recursion [TZ83]. Recursive [Bac82, Min85b, BPR82, BS87]. recursiveness [GJR83, GJ86]. reducible [Ros82, Sha87]. reducing [Car84]. reduction [HH86]. Reed [Tie85]. regions [AO81, CKSS81, CS84, Sak82]. regression [Mac81]. regular [Ros84, Tak81]. related [ES82]. relations [BPR82, MN84, SS85]. relationship [GH82]. Relationships [CK86]. relative [JS86]. relaxations [Bal85]. reliabilities [Hwa86]. reliability [BP82, DC81, SK86, SW87]. renewal [FS84]. representation [Bol85b, CNW80, DGK81a, Kir83, ST80, Toc84]. Representations [Pro82a, Bac82, CRS83, Doi87, SL83]. require [KKK83]. requirements [Car84]. requiring [KS81]. residual [Bha82]. residue [CW82]. resilience [Col87]. resource [Lan84]. restricted [BM87, Owe86]. result [Bol85a]. results [AJ81, DK87, NS81]. reversibility [Vic84]. reversible [HVW82]. Riccati [TS85, Wim82]. rigidity [LY82]. ring [Zem81]. Robinson [KV86]. Robinson-Schensted-Knuth [KV86]. robots [Tog86]. rock [DM86]. Room [DS87]. root [Gat83]. rounding [BT81]. rounds [HH82]. routing [WS82]. rule [Abe85]. Runs [SWG83].

S [CT15]. salesman [Cla84]. Sampling [Per81]. satellite [LL87]. satisfy [Her84a]. scalar [HMS87]. scaling [BPP81, DER85]. scaling-inequality [DER85]. Scheduling [DFL82, GJTY83, DW85, de 84a]. scheme [Ple86]. schemes [Bol85b, Had85, Per81]. Schensted [KV86, Vo81]. Schur [DGK81a, DGK83]. search [BM80, Boo82, Bos84, Cia85, MP84a]. second [Mer86, Tie85]. sections [GHW86]. selection [HT86]. self [MS81]. self-orthogonal [MS81]. selfadjoint [EGL87]. semi [Kri85, Lat82]. semi-iterative [Kri85]. semi-Markov [Lat82]. Semiantichains [WT81]. semidefinite [CDJ82, CDS84, HM84]. semigroup [AJ81]. semiorders [CR82, Doi87, Fis81]. Sensitive [Rot81]. Sensitivity [GM86, Elf81, Tog86]. Separating [FK84]. separator [GRE84]. Sequence [GHW86]. sequences [EHH87, FF80, Gam86, Gam87, HIS81, HG85]. sequential [BM80]. series [Col82, GJR83, HT86, SCC83]. series-parallel [HT86]. Set [KS81, CK87, DF83a, Gol82, Hoc83, HSY85, Sha87, VC82]. sets [BH85b, CH81, EZR85, FFF^{+}86, GP83, HHS82, JR85, Kan82b, Kan82a, Lih82, Pec81, Phi87, RS82, Wil86]. setups [GP83]. Shannon [Far86]. shape [MP84a, Rus80]. sharpening [Joh85]. shellable [BP82]. shifts [HMS87]. short [Gyö86]. shuffle [LLM84]. shuffle-exchange [LLM84]. side [GM87a]. sided [Ebe87]. sign [GH87a, HMS87]. signed [GLM83, GLM84]. Signs [TZ83]. similarity [ES82]. simple [AJ84, AT85, CNW80, CT82, Rab85, Tay82]. simplex [Neu83, Pro85]. Simplified [Hwa86]. Single [DW80, ST80]. Singular [CT15, Bar86, Buo86, BN87, CP83, Cam83, Cla86, Had85, Ros84, Tog86]. sitting [HS86]. Size [FK84, FW80]. skeletal [Sug83]. slides [SWG83]. small [GJ86, Hea87, Sax80]. small-bandwidth [Sax80]. smallest [Bar86]. Smith [KKS87]. soft [BS87]. sojourns [KW80b]. Solution [FS85, Szp87, BHP86, Cla86, EJ82, FW80, Gup86, Had85, Hal85, Hoc83, LM82, Min85b, OW85, Rab85, UP83, de 87]. solutions [BR82, BJW85, Buz82, LN80, Rot84, SS85, Vic85]. solvable [CP83]. solve [HS86]. solving [AG83, Gat83, SN85]. Some [AJ81, BR82, BPR82, DK87, GYY80, Her84b, KKP85, Sta80b, von81, Bha82, Bol85b, – Dok87, MN80, Per83, She80, Sug83, Szp87, Wan84, de 84b]. sonar [Gam87]. SOR [SN85]. Sorting [HH82]. space [Car84, CP86, CP87, GH87b]. spaces [ABBW87, Cha87]. spanning [Ber86, GH82]. Sparse [KOY86, CC86, GH87b, Liu87, NR83, Toc84]. spatially [BM80]. spatially-distributed [BM80]. specified [Gol80]. Spectra [Wan84, GH87a, Shr84]. spectral [BS86, BB87]. spectrum [Gup86]. Speed [Yao82, FB84]. Speed-up [Yao82]. Sperner [Pro82a, Sta80a]. spheroidal [Grü81]. splittings [Ros84]. Spreads [Kan82b, Kan82a]. square [GM87a]. squares [AO81, BCCL83, DGK80, DS87, Gil80a]. SS [LL87]. SS/TDMA [LL87]. SSOR [Kri85, Neu86]. Stability [Mat84, BH83, CN84, Wal85, Zem81]. stable [AO86, CDJ82, HHS82]. standby [DC81]. star [Tch85]. star-connected [Tch85]. starters [DS81]. Stationary [GM86, BP86, FP86, HR86, Hav86, Hey87, KMS84]. Steinberg [CBS85]. Steiner [FHM80, dPR82]. stochastic [ACN80, Smi81]. stock [Jon82]. stocks [Chu85a]. Stokes [Hal85]. storage [CRS83]. storing [Wal83]. straight [Car84]. straight-line [Car84]. stresses [WW83]. strings [CTPR85]. stripe [GS84]. striped [GS84]. Strong [FJ81, DS81]. strongly [BE84]. struction [de 84b]. structural [DG86, Ple86]. structure [GLL80, RS82, Wat87]. structures [ABBW87, BP85, CRS83, Sug83]. study [GGH80]. subadditive [BR82]. Subblock [Kir83]. subcodes [Tie85]. submatrices [BJOv87, KW80a]. suboptimal [SP84]. subset [Vo81]. subsets [DF83a, FFF^{+}86, Fis87, LR81]. subspaces [Bha83]. subsquare [GM87a]. Substitutes [Pro87]. subtraction [AMT86]. such [EJ82]. suitcase [HS86]. sum [DER85, LC84, Pro82b]. sums [Bar80, GH87a]. sun [CN84]. sun-free [CN84]. Super [BW86]. survey [RS85]. SVD [Shr84]. switch [DC81]. switches [Pip80]. switching [LL87]. Switchings [Tay82]. symmetric [CH87, DER85, Dia80, GJR83, HVW82, HMS87, KV86, Rab85, SL83]. symplectic [Bac82]. synopsis [SS85]. syntax [AB84]. system [AS80, Coo83, DFL82]. Systems [FK84, KOY86, BP82, Bou87, CM84, CP83, Cam83, Cla86, DC81, FFF^{+}86, Fis87, FP84, Had85, Hul84, Hwa86, MIN87, OW85, PR82, Pro83, Rot81, UP83, Wim84, dPR82, de 87].

tableaux [Gan81]. tandem [LN80]. TDMA [LL87]. technique [Gup86]. Telephone [BH86, HRS87]. telephones [Bum81]. Tensor [de 87]. Tensors [MW83]. terms [TZ83]. testing [CH80a, CH81, DK87, ES82, HHW81, HSD81, Hwa84, KCP82, Meh86]. their [Chu86, DGK83, EGL87, KSSS80, MW87, dN85]. them [Wal83]. theorem [Ber80b, CDJ82, Cha82, DF83b, FF80, GRE84, Gyö86, Pec81, Sta80a]. theorems [ANW84, BJW85, KKP85, Sta80b]. theoretic [GLM81]. theoretical [Ber80a]. theory [Chu84, EHH87, Fis83, GH85, JM87, Kör86, LM82]. those [Tre86]. Three [Hwa84, BS83, Bos84, Gam86]. three-valued [Gam86]. Threshold [CL84, Doi87, Fis83, HIS81, EOZ87, Gol82]. Threshold-bounded [Fis83]. ties [CK86]. tiesets [SCC83]. tight [BC81]. time [Cam83, Cla86, DFL82, FW80, Per87, Sax80]. time-band [Per87]. time-dependent [FW80]. time-varying [Cam83, Cla86]. titration [Shr84]. Toda [Chu84]. Toeplitz [Bul85, Grü81, Tre85, Tre86]. tool [Sta85]. Topological [MPS85, Chu85b]. total [LPHH84]. Totally [HKS85, Lew80, BL86, Coo83]. Totally-balanced [HKS85]. tournaments [Ber80a]. traceable [Fin85]. Traditional [KKK83]. transform [BS87, Per81]. transformation [Joh85]. transformations [BL83]. transforms [EH87, Mor86]. transitive [AJ84]. translation [Kan82b, Kan82a]. transmissions [Wes82]. transportation [AM81, BH85a]. travelling [Cla84]. tree [ACP87, CG82, Cha82, CDS83, CT80, Chu85b, GH82, JOv83, Liu87, MP84a, Wil86]. trees [AP86, Ber86, CH80b, CK87, DK87, FHM80, FP81, KS81, Len82, Moo83, PS83, Rus80]. trials [NS81]. triangular [HH86]. triangularization [MIN87]. triangulation [Lin87, Tod84b]. Triangulations [BM84, Sal84]. tridiagonal [CDJ82, Har86, Per87]. trinity [Ber80b]. triple [dPR82]. Tutte [Ber80b]. two [BW80, BFHMV84, Cam83, CH81, Chr83, CH80b, CGJ82, Dun80, GS84, HG85, SCC83]. two-chord [SCC83]. two-dimensional [CGJ82, HG85]. two-stripe [GS84]. type [Dun80, Lat82, Neu83].

ultrametric [ABBW87]. undiscounted [FS84]. unichain [WT81]. unimodality [CFG80]. unit [Rot84]. unitary [Bac82]. Unlabelled [CM84]. Updating [FP86]. Upper [Len82, BH85a]. urn [BP85, Gol85a]. use [Mol86]. Using [GM86, AG84, Gam87].

value [Bar86, Tog86]. valued [DGK81a, DGK81b, Gam86]. Values [Owe86, Bol87, GW80, Uhl85]. variable [BHP86, Hal85, MS86]. variable-complexity [MS86]. variables [Dun80, Sam80]. variational [GM87b]. varying [Cam83, Cla86]. Vector [CT15, KOY86, Phi87]. vectors [BN87, CW82, Phi87]. version [Grü81]. versions [Hwa84]. versus [FB84]. vertex [Gil80b, GJ86, PS86]. vertex-and [PS86]. Vertices [HHS82, BE84, Buc87, KS85, KT81]. via [HM84]. VLSI [CLR87, CGW87]. voltage [Ber83]. Volterra [Red86, RZ82, Red85a, Red85b]. Voronoı [CS84]. voting [Fis84].

walk [Mat87]. walks [KW80b, LT80]. watchmen [KKK83]. wave [Grü81]. weighing [CMW85]. Weight [Gag81, Gag83, MS81, Lin87]. Weighted [Mor86, AW80]. Weyl [Sta80a]. which [Her84a]. Whitney [IW81]. whose [Smi81]. width [Cha87, CGW87]. wiring [CGW87]. without [Wal83, Wes82, Wim82]. Worst [Won80]. Worst-case [Won80].

years [DF83b]. Young [Gan81].

zeros [BS86].

References

Albertson:1984:CAO

[AB84] Michael O. Albertson and David M. Berman. Combinatorial aspects of the orrery model of syntax. SIAM Journal on Algebraic and Discrete Methods, 5(3):451–456, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Aschbacher:1987:EUS

[ABBW87] Michael Aschbacher, Pierre Baldi, Eric B. Baum, and Richard M. Wilson. Embeddings of ultrametric spaces in finitedimensional structures. SIAM Journal on Algebraic and Discrete Methods, 8(4):564–577, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Abello:1985:ILM

[Abe85]  James M. Abello.  Intrin-

sic limitations of the majority rule, an algorithmic approach. SIAM Journal on Algebraic and Discrete Methods, 6 (1):133–144, January 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Aggarwal:1980:DSR

[ACN80] V. Aggarwal, R. Chandrasekaran, and K. P. K. Nair. Discounted stochastic ratio games. SIAM Journal on Algebraic and Discrete Methods, 1(2):201–210, June 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Arnborg:1987:CFE

[ACP87] Stefan Arnborg, Derek G.

Corneil, and Andrzej Proskurowski. Complexity of finding embeddings in a k-tree. SIAM Journal on Algebraic and Discrete Methods, 8(2):277–284, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Aulick:1983:IEE

[AG83] C. Mark Aulick and Thomas M.

Gallie. Isolating error effects in solving ill-posed problems. SIAM Journal on Algebraic and Discrete Methods, 4(3):371–376, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Aspvall:1984:GCU

[AG84] Bengt Aspvall and John R.

Gilbert. Graph coloring using eigenvalue decomposition. SIAM Journal on Algebraic and Discrete Methods, 5(4):526–538, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Araoz:1981:SRP

[AJ81] Julián Aráoz and Ellis L. Johnson. Some results on polyhedra of semigroup problems. SIAM Journal on Algebraic and Discrete Methods, 2(3):244– 258, September 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Abello:1984:HLT

[AJ84] James M. Abello and Charles R. Johnson. How large are transitive simple majority domains. SIAM Journal on Algebraic and Discrete Methods, 5(4):603–618, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Araoz:1985:MFN

[AJ85] Julián Aráoz and Ellis L. Johnson. Mappings and facets for nonabelian group problems. SIAM Journal on Algebraic and Discrete Methods, 6(2):171–188, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Aashtiani:1981:ECT

[AM81] H. Z. Aashtiani and T. L. Magnanti. Equilibria on a congested transportation network. SIAM Journal on Algebraic and Discrete Methods, 2(3):213–226, September 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Anderson:1987:CMH

[AMMT87] William N. Anderson, Jr., Michael E. Mays, Thomas D. Morley, and George E. Trapp. The contraharmonic mean of HSD matrices. SIAM Journal on Algebraic and Discrete Methods, 8(4):674–682, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Anderson:1986:CAS

[AMT86] W. N. Anderson, Jr., T. D. Morley, and G. E. Trapp. Cascade addition and subtraction of matrices. SIAM Journal on Algebraic and Discrete Methods, 7(4):609–626, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Aczel:1984:ATA

[ANW84] J. Aczél, C. T. Ng, and C. Wagner. Aggregation theorems for allocation problems. SIAM Journal on Algebraic and Discrete Methods, 5(1):1–8, March 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Albertson:1981:CRS

[AO81] Michael O. Albertson and Claire J. O’Keefe. Covering regions with squares. SIAM Journal on Algebraic and Discrete Methods, 2(3):240–243, September 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Ahac:1986:SMF

[AO86] Alan A. Ahac and D. D. Olesky.

A stable method for the LU factorization of M-matrices. SIAM Journal on Algebraic and Discrete Methods, 7(3):368–378, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Arnborg:1986:CRP

[AP86] Stefan Arnborg and Andrzej Proskurowski. Characterization and recognition of partial 3trees. SIAM Journal on Algebraic and Discrete Methods, 7(2):305–314, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Assmann:1981:BCH

[APSZ81] S. F. Assmann, G. W. Peck, M. M. Syslo, and J. Zak. The bandwidth of caterpillars with hairs of length 1 and 2*. SIAM Journal on Algebraic and Discrete Methods, 2(4):387– 393, December 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Arthurs:1980:COD

[AS80] E. Arthurs and B. W. Stuck. Controlling overload in a digital system. SIAM Journal on Algebraic and Discrete Methods, 1(2):232–250, June 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Alon:1985:CMS

[AT85] N. Alon and M. Tarsi. Covering multigraphs by simple circuits. SIAM Journal on Algebraic and Discrete Methods, 6(3):345–350, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Anderson:1986:IPM

[AT86] William N. Anderson, Jr. and George E. Trapp. Inverse problems for means of matrices. SIAM Journal on Algebraic and Discrete Methods, 7(2):188– 192, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Aczel:1980:CWA

[AW80] J. Aczél and C. Wagner. A characterization of weighted arithmetic means. SIAM Journal on Algebraic and Discrete Methods, 1(3):259–260, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Baclawski:1982:RAU

[Bac82] Kenneth Baclawski. Recursive algorithms for unitary and symplectic group representations. SIAM Journal on Algebraic and Discrete Methods, 3(4):592–605, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Balas:1985:DPH

[Bal85] Egon Balas. Disjunctive programming and a hierarchy of relaxations for discrete optimization problems. SIAM Journal on Algebraic and Discrete Methods, 6(3):466–486, July 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Baringhaus:1980:CDR

[Bar80] Ludwig Baringhaus. Characterization of distributions by random sums. SIAM Journal on Algebraic and Discrete Methods, 1(3):345–347, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Barnes:1982:APN

[Bar82] Earl R. Barnes. An algorithm for partitioning the nodes of a graph. SIAM Journal on Algebraic and Discrete Methods, 3(4):541–550, December 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Barnett:1984:MGP

[Bar84] Stephen Barnett. Multiplication of generalized polynomials, with applications to classical orthogonal polynomials. SIAM Journal on Algebraic and Discrete Methods, 5(4):457–462, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Barlow:1986:SPS

[Bar86] Jesse L. Barlow. On the smallest positive singular value of a singular M-matrix with applications to ergodic Markov chains. SIAM Journal on Algebraic and Discrete Methods, 7(3):414–424, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Barnett:1987:CBP

[Bar87] Stephen Barnett. Change of basis for products of orthogonal polynomials. SIAM Journal on Algebraic and Discrete Methods, 8(2):155–162, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Byers:1987:MMS

[BB87] W. Byers and A. Boyarsky.

Markov maps and the spectral radius of 0-1 matrices. SIAM Journal on Algebraic and Discrete Methods, 8(3):364–374, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Baker:1981:TAB

[BC81] B. S. Baker and E. G. Coffman, Jr. A tight asymptotic bound, for Next-Fit-Decreasing bin-packing. SIAM Journal on Algebraic and Discrete Methods, 2(2):147–152, June 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Baker:1983:AAM

[BCCL83] B. S. Baker, A. R. Calderbank, E. G. Coffman, Jr., and J. C. Lagarias. Approximation algorithms for maximizing the number of squares packed into a rectangle. SIAM Journal on Algebraic and Discrete Methods, 4(3):383–397, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Barak:1984:MNS

[BE84] Amnon B. Barak and Paul Erdős. On the maximal number of strongly independent vertices in a random acyclic directed graph. SIAM Journal on Algebraic and Discrete Methods, 5(4):508–514, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Berman:1980:GTA

[Ber80a] Kenneth A. Berman. A graphtheoretical approach to handicap ranking of tournaments and paired comparisons. SIAM Journal on Algebraic and Discrete Methods, 1(3):359–361, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Berman:1980:PTT

[Ber80b] Kenneth A. Berman. A proof of Tutte’s trinity theorem and a new determinant formula. SIAM Journal on Algebraic and Discrete Methods, 1(1):64–69, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Berman:1983:LEC

[Ber83] Kenneth A. Berman. Leaky electricity: 1-chain formulas for the current and voltage. SIAM Journal on Algebraic and Discrete Methods, 4(1):62–66, March 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Berman:1986:BST

[Ber86] Kenneth A. Berman. Bicycles and spanning trees. SIAM Journal on Algebraic and Discrete Methods, 7(1):1–12, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Berman:1987:EPD

[Ber87] Simeon M. Berman. An extension of Plackett’s differential equation for the multivariate normal density. SIAM Journal on Algebraic and Discrete Methods, 8(2):196–197, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Blake:1984:CLF

[BFHMV84] I. F. Blake, R. Fuji-Hara, R. C.

Mullin, and S. A. Vanstone. Computing logarithms in finite fields of characteristic two. SIAM Journal on Algebraic and Discrete Methods, 5(2):276–285, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Berman:1983:MDS

[BH83] Abraham Berman and Daniel Hershkowitz. Matrix diagonal stability and its implications. SIAM Journal on Algebraic and Discrete Methods, 4(3):377–382, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Barnes:1985:TPU

[BH85a] Earl R. Barnes and Alan J. Hoffman. On transportation problems with upper bounds on leading rectangles. SIAM Journal on Algebraic and Discrete Methods, 6(3):487–496, July 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Ben-Haim:1985:CSN

[BH85b] Yakov Ben-Haim. Convex sets and nondestructive assay. SIAM Journal on Algebraic and Discrete Methods, 6(4):688–706, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Bloom:1985:GDG

[BH85c] G. S. Bloom and D. F. Hsu. On graceful directed graphs. SIAM Journal on Algebraic and Discrete Methods, 6(3):519–536, July 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Berman:1986:TPF

[BH86] Kenneth A. Berman and Michael Hawrylycz. Telephone problems with failures. SIAM Journal on Algebraic and Discrete Methods, 7(1):13–17, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Bhattacharjee:1982:CMR

[Bha82] Manish C. Bhattacharjee. The class of mean residual lives and some consequences. SIAM Journal on Algebraic and Discrete Methods, 3(1):56–65, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Bhattacharyya:1983:GCI

[Bha83] S. P. Bhattacharyya. Generalized controllability, (A, B)- invariant subspaces and parameter invariant control. SIAM Journal on Algebraic and Discrete Methods, 4(4):529–533, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Burkardt:1986:DVM

[BHP86] J. Burkardt, C. Hall, and T. Porsching. The dual variable method for the solution of compressible fluid flow problems. SIAM Journal on Algebraic and Discrete Methods, 7(3):476–483, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Barrett:1987:IME

[BJOv87] Wayne W. Barrett, Charles R.

Johnson, D. D. Olesky, and P. van den Driessche. Inherited matrix entries: principal submatrices of the inverse. SIAM Journal on Algebraic and Discrete Methods, 8(3):313–322, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Butler:1985:NSQ

[BJW85] G. J. Butler, Charles R. Johnson, and H. Wolkowicz. Nonnegative solutions of a quadratic matrix equation arising from comparison theorems in ordinary differential equations. SIAM Journal on Algebraic and Discrete Methods, 6(1):47–53, January 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Brualdi:1983:LTN

[BL83] Richard A. Brualdi and Qiao Li. Linear transformations on nonnegative matrices preserving properties of irreducibility and full indecomposability. SIAM Journal on Algebraic and Discrete Methods, 4(1):116–131, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Broin:1986:DPA

[BL86] Martin W. Broin and Timothy J. Lowe. A dynamic programming algorithm for covering problems with (greedy) totally balanced constraint matrices. SIAM Journal on Algebraic and Discrete Methods, 7(3):348– 357, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Barker:1987:PEC

[BLP87] George Phillip Barker, Michael Laidacker, and George Poole. Projectionally exposed cones. SIAM Journal on Algebraic and Discrete Methods, 8(1):100–105, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Barnett:1980:SSS

[BM80] Arnold Barnett and John Mazzarino. On the sequential search for spatially-distributed events. SIAM Journal on Algebraic and Discrete Methods, 1(1):82–91, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Billera:1984:TOM

[BM84] Louis J. Billera and Beth Spellman Munson. Triangulations of oriented matroids and convex polytopes. SIAM Journal on Algebraic and Discrete Methods, 5(4):515–525, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Beker:1987:PRD

[BM87] Henry Beker and Chris Mitchell.

Permutations with restricted displacement. SIAM Journal on Algebraic and Discrete Methods, 8(3):338–363, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Byers:1987:SVL

[BN87] Ralph Byers and Stephen Nash.

On the singular “vectors” of the Lyapunov operator. SIAM Journal on Algebraic and Discrete Methods, 8(1):59–66, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic). See note [CT15].

Bolger:1983:BIM

[Bol83] Edward M. Bolger. The Banzhaf index for multicandidate presidential elections. SIAM Journal on Algebraic and Discrete Methods, 4(4):442–458, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Boley:1985:PRL

[Bol85a] Daniel Boley. A perturbation result for linear control problems. SIAM Journal on Algebraic and Discrete Methods, 6 (1):66–72, January 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Bolger:1985:MOP

[Bol85b] Edward M. Bolger. Monotonicity and other paradoxes in some proportional representation schemes. SIAM Journal on Algebraic and Discrete Methods, 6(2):283–291, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Bolger:1987:CEV

[Bol87] Edward M. Bolger. A class of efficient values for games in partition function form. SIAM Journal on Algebraic and Discrete Methods, 8(3):460–466, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Booth:1982:OOS

[Boo82] R. S. Booth. An odd order search problem. SIAM Journal on Algebraic and Discrete Methods, 3 (1):135–143, March 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Bostock:1984:DSP

[Bos84] F. A. Bostock. On a discrete search problem on three arcs. SIAM Journal on Algebraic and Discrete Methods, 5(1):94– 100, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Boucher:1984:HCA

[Bou84] Andy Boucher. It’s hard to color antirectangles. SIAM Journal on Algebraic and Discrete Methods, 5(2):162–163, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Bouchet:1987:DDE

[Bou87] André Bouchet. Digraph decompositions and Eulerian systems. SIAM Journal on Algebraic and Discrete Methods, 8(3):323–337, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Ball:1982:BRP

[BP82] Michael O. Ball and J. Scott Provan. Bounds on the reliability polynomial for shellable independence systems. SIAM Journal on Algebraic and Discrete Methods, 3(2):166–181, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Bagchi:1985:ANG

[BP85] A. Bagchi and A. K. Pal. Asymptotic normality in the generalized Polya-Eggenberger urn model, with an application to computer data structures. SIAM Journal on Algebraic and Discrete Methods, 6(3):394–405, July 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Barker:1986:CIC

[BP86] G. P. Barker and R. J. Plemmons. Convergent iterations for computing stationary distributions of Markov chains. SIAM Journal on Algebraic and Discrete Methods, 7(3):390–398, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Berman:1981:DSO

[BPP81] A. Berman, B. N. Parlett, and R. J. Plemmons. Diagonal scaling to an orthogonal matrix. SIAM Journal on Algebraic and Discrete Methods, 2 (1):57–65, March 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Batty:1982:SRR

[BPR82] C. J. K. Batty, M. J. Pelling, and D. G. Rogers. Some recurrence relations of recursive minimization. SIAM Journal on Algebraic and Discrete Methods, 3 (1):13–29, March 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Batty:1982:SMS

[BR82] C. J. K. Batty and D. G. Rogers. Some maximal solutions of the generalized subadditive inequality. SIAM Journal on Algebraic and Discrete Methods, 3(3):369– 378, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Brawley:1982:CMD

[Bra82] J. V. Brawley. Counting matrices by Drazin index. SIAM Journal on Algebraic and Discrete Methods, 3(1):30–34, March 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Barnes:1983:OLQ

[BS83] E. S. Barnes and N. J. A. Sloane.

The optimal lattice quantizer in three dimensions. SIAM Journal on Algebraic and Discrete Methods, 4(1):30–41, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Brualdi:1986:SRC

[BS86] Richard A. Brualdi and Ernie S.

Solheid. On the spectral radius of complementary acyclic matrices of zeros and ones. SIAM Journal on Algebraic and Discrete Methods, 7(2):265–272, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Beery:1987:RHT

[BS87] Yair Be’ery and Jakov Snyders.

A recursive Hadamard transform optimal soft decision decoding algorithm. SIAM Journal on Algebraic and Discrete Methods, 8(4):778–789, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Baum:1981:IRP

[BT81] S. Baum and L. E. Trotter, Jr. Integer rounding for polymatroid and branching optimization problems. SIAM Journal on Algebraic and Discrete Methods, 2(4):416–425, December 1981. CODEN SJAMDU.

ISSN 0196-5212 (print), 2168-345X (electronic).

Buck:1986:ED

[Buc86] Marshall W. Buck. Expanders and diffusers. SIAM Journal on Algebraic and Discrete Methods, 7(2):282–304, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Buchta:1987:NVR

[Buc87]  Christian Buchta.  On the

number of vertices of random polyhedra with a given number of facets. SIAM Journal on Algebraic and Discrete Methods, 8(1):85–92, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Bultheel:1985:ABT

[Bul85] Adhemar Bultheel. The asymptotic behavior of Toeplitz determinants generated by the Laurent coefficients of a meromorphic function. SIAM Journal on Algebraic and Discrete Methods, 6(4):624–629, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Bumby:1981:PT

[Bum81] Richard T. Bumby. A problem with telephones. SIAM Journal on Algebraic and Discrete Methods, 2(1):13–18, March 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Buoni:1986:IFS

[Buo86] J. J. Buoni. Incomplete factorization of singular M-matrices. SIAM Journal on Algebraic and Discrete Methods, 7(2):193–198, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Buzytsky:1982:EFN

[Buz82] P. L. Buzytsky. An effective formula for the number of solutions of linear Boolean equations. SIAM Journal on Algebraic and Discrete Methods, 3(2):182–186, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Bertram:1980:DPT

[BW80] Edward A. Bertram and Victor K. Wei. Decomposing a permutation into two large cycles: an enumeration. SIAM Journal on Algebraic and Discrete Methods, 1(4):450–461, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Boesch:1986:SLC

[BW86] F. T. Boesch and J. F. Wang. Super line-connectivity properties of circulant graphs. SIAM Journal on Algebraic and Discrete Methods, 7(1):89–98, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Campbell:1983:ITL

[Cam83] Stephen L. Campbell. Index two linear time-varying singular systems of differential equations. SIAM Journal on Algebraic and Discrete Methods, 4(2):237–243, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Carlson:1984:RSR

[Car84] David A. Carlson. On reducing the space requirements of a straight-line algorithm. SIAM Journal on Algebraic and Discrete Methods, 5(3):442–450, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Carter:1985:IPA

[CBS85] H. W. Carter, M. A. Breuer, and Z. A. Syed. Incremental processing applied to Munkres’ algorithm and its application in Steinberg’s placement procedure. SIAM Journal on Algebraic and Discrete Methods, 6(2):210–219, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Coleman:1986:CCP

[CC86] Thomas F. Coleman and Jin-Yi Cai. The cyclic coloring problem and estimation of sparse Hessian matrices. SIAM Journal on Algebraic and Discrete Methods, 7(2):221–235, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Colbourn:1982:CBD

[CCPR82] Charles J. Colbourn, Marlene J. Colbourn, Kevin T. Phelps, and Vojtěch Rödl. Coloring block designs is NP-complete.

SIAM Journal on Algebraic and Discrete Methods, 3(3):305–307, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Carlson:1982:SLT

[CDJ82] David Carlson, B. N. Datta, and Charles R. Johnson. A semidefinite Lyapunov theorem and the characterization of tridiagonal D-stable matrices. SIAM Journal on Algebraic and Discrete Methods, 3(3):293–304, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chaiken:1983:ODT

[CDS83] S. Chaiken, A. K. Dewdney, and P. J. Slater. An optimal diagonal tree code. SIAM Journal on Algebraic and Discrete Methods, 4(1):42–49, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Carlson:1984:CMP

[CDS84] David Carlson, B. N. Datta, and Hans Schneider. On the controllability of matrix pairs (A, K) with K positive semidefinite. SIAM Journal on Algebraic and Discrete Methods, 5(3):346– 350, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chung:1980:ULE

[CFG80] F. R. K. Chung, P. C. Fishburn, and R. L. Graham. On unimodality for linear extensions of partial orders. SIAM Journal on Algebraic and Discrete Methods, 1(4):405–410, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Camerini:1982:BPT

[CG82] Paolo M. Camerini and Giulia Galbiati. The bounded path tree problem. SIAM Journal on Algebraic and Discrete Methods, 3(4):474–484, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chung:1982:PTD

[CGJ82] F. R. K. Chung, M. R. Garey, and D. S. Johnson. On packing two-dimensional bins. SIAM Journal on Algebraic and Discrete Methods, 3(1):66–76, March 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Coppersmith:1987:DMC

[CGW87] D. Coppersmith, I. Gopal, and C. K. Wong. Distribution of the minimum channel width in VLSI wiring. SIAM Journal on Algebraic and Discrete Methods, 8(2):285–290, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chang:1980:GTP

[CH80a] Gerard J. Chang and F. K. Hwang. A group testing problem. SIAM Journal on Algebraic and Discrete Methods, 1(1):21– 24, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chung:1980:CPT

[CH80b] F. R. K. Chung and F. K.

Hwang. The connection patterns of two complete binary trees. SIAM Journal on Algebraic and Discrete Methods, 1(3):322–335, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chang:1981:GTP

[CH81] Gerard J. Chang and F. K.

Hwang. A group testing problem on two disjoint sets. SIAM Journal on Algebraic and Discrete Methods, 2(1):35–38, March 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Choudhury:1987:COS

[CH87] Dipa Choudhury and Roger A.

Horn. A complex orthogonalsymmetric analog of the polar decomposition. SIAM Journal on Algebraic and Discrete Methods, 8(2):219–225, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chaiken:1982:CPA

[Cha82] Seth Chaiken. A combinatorial proof of the all minors matrix tree theorem. SIAM Journal on Algebraic and Discrete Methods, 3(3):319–329, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chaiken:1983:MAB

[Cha83] Seth Chaiken. A matroid abstraction of the Bott–Duffin constrained inverse. SIAM Journal on Algebraic and Discrete Methods, 4(4):467–475, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chaiken:1987:ELW

[Cha87] Seth Chaiken. Extremal length and width of blocking polyhedra, Kirchhoff spaces and multiport networks. SIAM Journal on Algebraic and Discrete Methods, 8(4):635–645, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Christen:1983:ODT

[Chr83] C. Christen. Optimal detection of two complementary defectives. SIAM Journal on Algebraic and Discrete Methods, 4 (1):101–110, March 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chung:1981:DG

[Chu81] F. R. K. Chung. On the decomposition of graphs. SIAM Journal on Algebraic and Discrete Methods, 2(1):1–12, March 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chu:1984:GTF

[Chu84] Moody T. Chu. The generalized Toda flow, the QR algorithm and the center manifold theory. SIAM Journal on Algebraic and Discrete Methods, 5(2):187– 201, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chung:1985:NNI

[Chu85a] Chia-Shin Chung. A note on the number of invariant capital stocks. SIAM Journal on Algebraic and Discrete Methods, 6(2):242–244, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chung:1985:CTB

[Chu85b] Fan R. K. Chung. On the cutwidth and the topological bandwidth of a tree. SIAM Journal on Algebraic and Discrete Methods, 6(2):268–277, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chu:1986:CLE

[Chu86] Moody T. Chu. Curves on S^{n−1} that lead to eigenvalues or their means of a matrix. SIAM Journal on Algebraic and Discrete Methods, 7(3):425– 432, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Ciancutti:1985:DSM

[Cia85] Mark Ciancutti. On discrete search for a multiple number of objects. SIAM Journal on Algebraic and Discrete Methods, 6(2):335–340, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Cook:1986:RBM

[CK86] Wade D. Cook and Moshe Kress.

Relationships between l^{1} metrics on rankings: the case of ties. SIAM Journal on Algebraic and Discrete Methods, 7(3):445– 451, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Corneil:1987:DPA

[CK87] D. G. Corneil and J. M. Keil. A dynamic programming approach to the dominating set problem on k-trees. SIAM Journal on Algebraic and Discrete Methods, 8(4):535–543, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chaiken:1981:CRR

[CKSS81] Seth Chaiken, Daniel J. Kleitman, Michael Saks, and James Shearer. Covering regions by rectangles. SIAM Journal on Algebraic and Discrete Methods, 2(4):394–410, December 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Cozzens:1984:TDG

[CL84] Margaret B. Cozzens and Rochelle Leibowitz. Threshold dimension of graphs. SIAM Journal on Algebraic and Discrete Methods, 5(4):579–595, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Claus:1984:NFT

[Cla84] A. Claus. A new formulation for the travelling salesman problem. SIAM Journal on Algebraic and Discrete Methods, 5 (1):21–25, March 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Clark:1986:DMN

[Cla86] Kenneth D. Clark. Difference methods for the numerical solution of time-varying singular systems of differential equations. SIAM Journal on Algebraic and Discrete Methods, 7(2):236–246, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chung:1987:EGB

[CLR87] Fan R. K. Chung, Frank Thomson Leighton, and Arnold L. Rosenberg. Embedding graphs in books: a layout problem with applications to VLSI design. SIAM Journal on Algebraic and Discrete Methods, 8(1):33– 58, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Camerini:1984:UPS

[CM84] P. M. Camerini and F. Maffioli. Unlabelled partition systems: optimization and complexity. SIAM Journal on Algebraic and Discrete Methods, 5(3):426–441, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Cheng:1985:OWD

[CMW85] Ching-Shui Cheng, Joseph C.

Masaro, and Chi Song Wong. Optimal weighing designs. SIAM Journal on Algebraic and Discrete Methods, 6(2):259–267, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chang:1984:DSP

[CN84] Gerard J. Chang and George L. Nemhauser. The k-domination and k-stability problems on sunfree chordal graphs. SIAM Journal on Algebraic and Discrete Methods, 5(3):332–345, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chang:1985:CPG

[CN85] Gerard J. Chang and George L. Nemhauser. Covering, packing and generalized perfection. SIAM Journal on Algebraic and Discrete Methods, 6(1):109– 132, January 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Cornuejols:1980:CRS

[CNW80] Gerard Cornuejols, George L. Nemhauser, and Laurence A. Wolsey. A canonical representation of simple plant location problems and its applications. SIAM Journal on Algebraic and Discrete Methods, 1(3):261–272, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Colbourn:1982:FSM

[Col82] Charles J. Colbourn. Farey series and maximal outerplanar graphs. SIAM Journal on Algebraic and Discrete Methods, 3(2):187–189, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Colbourn:1987:NR

[Col87] Charles J. Colbourn. Network resilience. SIAM Journal on Algebraic and Discrete Methods, 8(3):404–409, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Cook:1983:MTD

[Coo83] William Cook. A minimal totally dual integral defining system for the b-matching polyhedron. SIAM Journal on Algebraic and Discrete Methods, 4(2):212–220, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Campbell:1983:CFS

[CP83] Stephen L. Campbell and Linda R. Petzold. Canonical forms and solvable singular systems of differential equations. SIAM Journal on Algebraic and Discrete Methods, 4(4):517–521, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Coleman:1986:NSP

[CP86] Thomas F. Coleman and Alex Pothen. The null space problem. I. Complexity. SIAM Journal on Algebraic and Discrete Methods, 7(4):527–537, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Coleman:1987:NSP

[CP87] Thomas F. Coleman and Alex Pothen. The null space problem. II. Algorithms. SIAM Journal on Algebraic and Discrete Methods, 8(4):544–563, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Cozzens:1982:DSD

[CR82] Margaret B. Cozzens and Fred S. Roberts. Double semiorders and double indifference graphs. SIAM Journal on Algebraic and Discrete Methods, 3(4):566–583, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chung:1983:PSR

[CRS83] F. R. K. Chung, A. L. Rosenberg, and Lawrence Snyder. Perfect storage representations for families of data structures. SIAM Journal on Algebraic and Discrete Methods, 4(4):548– 565, December 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Conway:1984:VRC

[CS84] J. H. Conway and N. J. A. Sloane. On the Voronoĭ regions of certain lattices. SIAM Journal on Algebraic and Discrete Methods, 5(3):294–305, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chhetry:1987:PDB

[CS87] Devendra Chhetry and Allan R. Sampson. A projection decomposition for bivariate discrete probability distributions. SIAM Journal on Algebraic and Discrete Methods, 8(3):501–509, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chandrasekaran:1980:ACC

[CT80] R. Chandrasekaran and A. Tamir.

An O((n log p)^{2}) algorithm for the continuous p-center problem on a tree. SIAM Journal on Algebraic and Discrete Methods, 1(4):370–375, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Cornuejols:1982:PAS

[CT82] Gerard Cornuejols and Jean-Michel Thizy. A primal approach to the simple plant location problem. SIAM Journal on Algebraic and Discrete Methods, 3(4):504–510, December 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Chen:2015:NSV

[CT15] Sheng Chen and Yunbo Tian.

Note on “On the Singular ‘Vectors’ of the Lyapunov Operator” by R. Byers and S. Nash. SIAM Journal on Matrix Analysis and Applications, 36(3): 1069–1072, ???? 2015. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). See [BN87].

Chung:1985:CSP

[CTPR85] F. R. K. Chung, R. E. Tarjan, W. J. Paul, and R. Reischuk. Coding strings by pairs of strings. SIAM Journal on Algebraic and Discrete Methods, 6(3):445–461, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Culik:1984:MM

[Čul84] Karel Čulík. On metapaths in metagraphs. SIAM Journal on Algebraic and Discrete Methods, 5(4):551–568, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Cunningham:1982:DDG

[Cun82] William H. Cunningham. Decomposition of directed graphs. SIAM Journal on Algebraic and Discrete Methods, 3(2):214–228, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Calderbank:1982:MVB

[CW82] Robert Calderbank and David B. Wales. Multiplying vectors in binary quadratic residue codes. SIAM Journal on Algebraic and Discrete Methods, 3(1):43–55, March 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Cozzens:1987:MCE

[CW87] Margaret B. Cozzens and Shu Shih Y. Wu. On minimum critically n-edge-connected graphs. SIAM Journal on Algebraic and Discrete Methods, 8(4):659–669, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Davis:1986:CFP

[Dav86] George J. Davis. Column LU factorization with pivoting on a message-passing multiprocessor. SIAM Journal on Algebraic and Discrete Methods, 7(4):538– 550, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Downs:1981:RSS

[DC81] T. Downs and P. K. W. Chan.

The reliability of standby systems with a faulty switch. SIAM Journal on Algebraic and Discrete Methods, 2(4):464– 471, December 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Daykin:1985:OPM

[DD85] David E. Daykin and Jacqueline W. Daykin. Order preserving maps and linear extensions of a finite poset. SIAM Journal on Algebraic and Discrete Methods, 6(4):738–748, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

deWerra:1984:PSL

[de 84a]  D. de Werra.  Preemptive

scheduling, linear programming and network flows. SIAM Journal on Algebraic and Discrete Methods, 5(1):11–20, March 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

deWerra:1984:SPS

[de 84b] Dominique de Werra. On some properties of the struction of a graph. SIAM Journal on Algebraic and Discrete Methods, 5(2):239–243, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

dePillis:1987:TES

[de 87] John de Pillis. Tensor equivalents for solution of linear systems: a parallel algorithm. SIAM Journal on Algebraic and Discrete Methods, 8(3):291–312, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Demetrovics:1980:CKA

[Dem80] J. Demetrovics. Candidate keys and antichains. SIAM Journal on Algebraic and Discrete Methods, 1(1):92, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Dantzig:1985:DSI

[DER85] George B. Dantzig, B. Curtis Eaves, and Uriel G. Rothblum. A decomposition and scaling-inequality for line-sum-symmetric nonnegative matrices. SIAM Journal on Algebraic and Discrete Methods, 6(2):237– 241, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Daykin:1983:ISS

[DF83a] David E. Daykin and Peter Frankl. Inequalities for subsets of a set and KLYM posets. SIAM Journal on Algebraic and Discrete Methods, 4(1):67–69, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Deza:1983:EKR

[DF83b] M. Deza and P. Frankl. Erdős-Ko-Rado theorem—22 years later. SIAM Journal on Algebraic and Discrete Methods, 4(4):419–431, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Deuermeyer:1982:SMM

[DFL82] Bryan L. Deuermeyer, Donald K. Friesen, and Michael A. Langston. Scheduling to maximize the minimum processor finish time in a multiprocessor system. SIAM Journal on Algebraic and Discrete Methods, 3(2):190– 196, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Duff:1986:CSI

[DG86] I. S. Duff and C. W. Gear.

Computing the structural index. SIAM Journal on Algebraic and Discrete Methods, 7(4):594– 603, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Delsarte:1980:PLS

[DGK80] Ph. Delsarte, Y. Genin, and Y. Kamp. Planar least-squares inverse polynomials. II. Asymptotic behavior. SIAM Journal on Algebraic and Discrete Methods, 1(3):336–344, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Delsarte:1981:GSR

[DGK81a] P. Delsarte, Y. Genin, and Y. Kamp. Generalized Schur representation of matrix-valued functions. SIAM Journal on Algebraic and Discrete Methods, 2(2):94–107, 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Delsarte:1981:HPM

[DGK81b] Ph. Delsarte, Y. Genin, and Y. Kamp. Half-plane minimization of matrix-valued quadratic functionals. SIAM Journal on Algebraic and Discrete Methods, 2(2):192–211, June 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Delsarte:1983:ECH

[DGK83] Philippe Delsarte, Y. Genin, and Y. Kamp. Equivalence classes of Hermitian matrices and their Schur parametrization. SIAM Journal on Algebraic and Discrete Methods, 4(3):279– 289, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Du:1982:MCF

[DH82] Ding Zhu Du and F. K. Hwang. Minimizing a combinatorial function. SIAM Journal on Algebraic and Discrete Methods, 3(4):523–528, December 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Du:1987:OAC

[DH87] D. Z. Du and F. K. Hwang. Optimal assignments for consecutive-2 graphs. SIAM Journal on Algebraic and Discrete Methods, 8(3):510–518, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Diamond:1980:AEC

[Dia80] Harvey Diamond. Asymptotic equilibria in a class of N-person symmetric games. SIAM Journal on Algebraic and Discrete Methods, 1(1):121–136, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Djidjev:1982:PPP

[Dji82] Hristo Nicolov Djidjev. On the problem of partitioning planar graphs. SIAM Journal on Algebraic and Discrete Methods, 3(2):229–240, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

deBruijn:1982:MAC

[dK82] N. G. de Bruijn and D. A.

Klarner. Multisets of aperiodic cycles. SIAM Journal on Algebraic and Discrete Methods, 3(3):359–368, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Du:1987:SCR

[DK87] Ding Zhu Du and Ker-I Ko. Some completeness results on decision trees and group testing. SIAM Journal on Algebraic and Discrete Methods, 8(4):762– 777, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Delport:1986:MMI

[DM86] J. L. Delport and D. H. Martin. A multiplier method for identifying keyblocks in excavations through jointed rock. SIAM Journal on Algebraic and Discrete Methods, 7(2):321–330, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

dePillis:1985:EPH

[dN85] John de Pillis and Michael Neumann. The effect of the perturbation of Hermitian matrices on their eigenvectors. SIAM Journal on Algebraic and Discrete Methods, 6(2):201–209, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Doignon:1987:TRM

[Doi87] Jean-Paul Doignon. Threshold representations of multiple semiorders. SIAM Journal on Algebraic and Discrete Methods, 8(1):77–84, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

deBrandes:1982:CST

[dPR82] Marcia de Brandes, Kevin T.

Phelps, and Vojtěch Rödl. Coloring Steiner triple systems. SIAM Journal on Algebraic and Discrete Methods, 3(2):241–249, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Drews:1982:PAD

[Dre82] W. P. Drews. Practical applications of discrete mathematical programming in Exxon. SIAM Journal on Algebraic and Discrete Methods, 3(4):453– 464, December 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Dinitz:1981:FAF

[DS81] J. H. Dinitz and D. R. Stinson. A fast algorithm for finding strong starters. SIAM Journal on Algebraic and Discrete Methods, 2 (1):50–56, March 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Dinitz:1987:HCA

[DS87] J. H. Dinitz and D. R. Stinson. A hill-climbing algorithm for the construction of one-factorizations and Room squares. SIAM Journal on Algebraic and Discrete Methods, 8(3):430–438, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Dunkl:1980:OPT

[Dun80] Charles F. Dunkl. Orthogonal polynomials in two variables of q-Hahn and q-Jacobi type. SIAM Journal on Algebraic and Discrete Methods, 1(2):137–151, June 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Drezner:1980:SFD

[DW80] Z. Drezner and G. O. Wesolowsky. Single facility l_{p}-distance minimax location. SIAM Journal on Algebraic and Discrete Methods, 1(3):315–321, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Dolev:1985:PSO

[DW85] Danny Dolev and Manfred K. Warmuth. Profile scheduling of opposing forests and level orders. SIAM Journal on Algebraic and Discrete Methods, 6(4):665– 687, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Eberlein:1987:OSJ

[Ebe87] P. J. Eberlein. On one-sided Jacobi methods for parallel computation. SIAM Journal on Algebraic and Discrete Methods, 8(4):790–796, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Ellis:1987:ISE

[EGL87] Robert L. Ellis, Israel Gohberg, and David C. Lay. Invertible selfadjoint extensions of band matrices and their entropy. SIAM Journal on Algebraic and Discrete Methods, 8(3):483–500, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Eberly:1987:GCM

[EH87] David Eberly and Paul Hartung.

Group convolutions and matrix transforms. SIAM Journal on Algebraic and Discrete Methods, 8(2):263–276, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Ehrenfeucht:1987:QST

[EHH87] Andrzej Ehrenfeucht, Jeffrey Haemer, and David Haussler. Quasimonotonic sequences: theory, algorithms and applications. SIAM Journal on Algebraic and Discrete Methods, 8(3):410–429, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Egawa:1982:NMS

[EJ82] Yoshimi Egawa and S. K. Jain.

Nonnegative matrices A such that Ax = b has nonnegative best approximate solution. SIAM Journal on Algebraic and Discrete Methods, 3(2):197– 213, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Elfving:1981:SGM

[Elf81] Tommy Elfving. On the sensitivity of the gravity model. SIAM Journal on Algebraic and Discrete Methods, 2(1):19–24, March 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Erdos:1987:BTD

[EOZ87] Paul Erdős, Edward T. Ordman, and Yechezkel Zalcstein. Bounds on threshold dimension and disjoint threshold coverings. SIAM Journal on Algebraic and Discrete Methods, 8(2):151– 154, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Engel:1982:ATD

[ES82] Gernot M. Engel and Hans Schneider. Algorithms for testing the diagonal similarity of matrices and related problems. SIAM Journal on Algebraic and Discrete Methods, 3(4):429–438, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

El-Zahar:1985:EJC

[EZR85] M. H. El-Zahar and I. Rival. Examples of jump-critical ordered sets. SIAM Journal on Algebraic and Discrete Methods, 6(4):713– 720, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Farber:1986:ASC

[Far86] Martin Farber. An analogue of the Shannon capacity of a graph. SIAM Journal on Algebraic and Discrete Methods, 7(1):67–72, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Friedman:1984:EVS

[FB84] Nathan Friedman and Abraham Boyarsky. Entropy versus speed in ergodic Markov maps. SIAM Journal on Algebraic and Discrete Methods, 5(1):82–93, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Frankl:1980:EKR

[FF80] Peter Frankl and Zoltán Füredi.

The Erdős-Ko-Rado theorem for integer sequences. SIAM Journal on Algebraic and Discrete Methods, 1(4):376–381, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Fishburn:1986:PIS

[FFF^{+}86] P. C. Fishburn, P. Frankl, D. Freed, J. C. Lagarias, and A. M. Odlyzko. Probabilities for intersecting systems and random subsets of finite sets. SIAM Journal on Algebraic and Discrete Methods, 7(1):73–79, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Farley:1980:RST

[FHM80] Arthur M. Farley, Stephen T.

Hedetniemi, and Sandra L. Mitchell. Rectilinear Steiner trees in rectangle trees. SIAM Journal on Algebraic and Discrete Methods, 1(1):70–81, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Fink:1985:RNT

[Fin85] John Frederick Fink. Randomly near-traceable graphs. SIAM Journal on Algebraic and Discrete Methods, 6(2):251–258, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Fishburn:1981:MSI

[Fis81]  Peter C. Fishburn.  Maxi-

mum semiorders in interval orders. SIAM Journal on Algebraic and Discrete Methods, 2(2):127– 135, 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Fishburn:1983:TBI

[Fis83] Peter C. Fishburn. Thresholdbounded interval orders and a theory of picycles. SIAM Journal on Algebraic and Discrete Methods, 4(3):290–305, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Fishburn:1984:DMV

[Fis84] Peter C. Fishburn. Discrete mathematics in voting and group choice. SIAM Journal on Algebraic and Discrete Methods, 5 (2):263–275, June 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Fishburn:1987:LSD

[Fis87] P. C. Fishburn. Latent subsets for dual intersecting systems. SIAM Journal on Algebraic and Discrete Methods, 8(4):628–634, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Flinchbaugh:1981:SCD

[FJ81] B. E. Flinchbaugh and L. K.

Jones. Strong connectivity in directional nearest-neighbor graphs. SIAM Journal on Algebraic and Discrete Methods, 2(4):461–463, December 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Farber:1986:CGH

[FJ86] Martin Farber and Robert E.

Jamison. Convexity in graphs and hypergraphs. SIAM Journal on Algebraic and Discrete Methods, 7(3):433–444, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Fredman:1984:SSS

[FK84] Michael L. Fredman and János Komlós. On the size of separating systems and families of perfect hash functions. SIAM Journal on Algebraic and Discrete Methods, 5(1):61–68, March 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Farley:1981:BTM

[FP81] Arthur M. Farley and Andrzej Proskurowski. Broadcasting in trees with multipole originators. SIAM Journal on Algebraic and Discrete Methods, 2(4):381–386, December 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Funderlic:1984:CDI

[FP84] R. E. Funderlic and R. J. Plemmons. A combined direct-iterative method for certain M-matrix linear systems. SIAM Journal on Algebraic and Discrete Methods, 5(1):33–42, March 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Funderlic:1986:UFC

[FP86] R. E. Funderlic and R. J. Plemmons. Updating LU factorizations for computing stationary distributions. SIAM Journal on Algebraic and Discrete Methods, 7(1):30–42, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Friedland:1980:GPN

[FS80] Shmuel Friedland and Hans Schneider. The growth of powers of a nonnegative matrix. SIAM Journal on Algebraic and Discrete Methods, 1(2):185–200, June 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Federgrun:1984:FPA

[FS84] A. Federgrün and P. J. Schweitzer. A fixed point approach to undiscounted Markov renewal programs. SIAM Journal on Algebraic and Discrete Methods, 5(4):539–550, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Fu:1985:SDL

[FS85] Sheau Wei Fu and Mahmoud E.

Sawan. Solution of the discrete Lyapunov equation. SIAM Journal on Algebraic and Discrete Methods, 6(2):341–344, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Friedman:1980:ASI

[FW80] Moshe Friedman and Jeffrey L.

Winter. An asymptotic solution of inventory lot-size models with homogeneous timedependent demand functions. SIAM Journal on Algebraic and Discrete Methods, 1(3):300–314, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Fisher:1982:GHC

[FW82] Marshall L. Fisher and Laurence A. Wolsey. On the greedy heuristic for continuous covering and packing problems. SIAM Journal on Algebraic and Discrete Methods, 3(4):584–591, December 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Gagola:1981:WEN

[Gag81] Stephen M. Gagola, Jr. Weight enumerators of normalized codes. SIAM Journal on Algebraic and Discrete Methods, 2(4):347– 380, December 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Gagola:1983:WEN

[Gag83] Stephen M. Gagola, Jr. Weight enumerators of normalized codes. II. The Hermitian case. SIAM Journal on Algebraic and Discrete Methods, 4(3):317–354, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Galambos:1986:PLB

[Gal86] G. Galambos. Parametric lower bound for on-line bin-packing. SIAM Journal on Algebraic and Discrete Methods, 7(3):362–367, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Games:1986:GST

[Gam86] Richard A. Games. The geometry of m-sequences: three-valued crosscorrelations and quadrics in finite projective geometry. SIAM Journal on Algebraic and Discrete Methods, 7(1):43–52, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Games:1987:ACS

[Gam87] Richard A. Games. An algebraic construction of sonar sequences using M-sequences. SIAM Journal on Algebraic and Discrete Methods, 8(4):753–761, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Gansner:1981:ADY

[Gan81] Emden R. Gansner. Acyclic digraphs, Young tableaux and nilpotent matrices. SIAM Journal on Algebraic and Discrete Methods, 2(4):429–440, December 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Gati:1983:CSP

[Gat83] Georg Gati. The complexity of solving polynomial equations by prime root extractions. SIAM Journal on Algebraic and Discrete Methods, 4(1):70–71, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Greenberg:1980:CPA

[GGH80] James Greenberg, Curtis Greene, and Stuart Hastings. A combinatorial problem arising in the study of reaction-diffusion equations. SIAM Journal on Algebraic and Discrete Methods, 1(1):34–42, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Granot:1982:RBC

[GH82] Daniel Granot and Gur Huberman. The relationship between convex games and minimum cost spanning tree games: a case for permutationally convex games. SIAM Journal on Algebraic and Discrete Methods, 3(3):288–292, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Gustafson:1985:GTF

[GH85] Karl Gustafson and Robert Hartman. Graph theory and fluid dynamics. SIAM Journal on Algebraic and Discrete Methods, 6(4):643–656, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Garloff:1987:SMH

[GH87a] Jürgen Garloff and Volker Hattenbach. The spectra of matrices having sums of principal minors with alternating sign. SIAM Journal on Algebraic and Discrete Methods, 8(1):106–107, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Gilbert:1987:CSB

[GH87b] John R. Gilbert and Michael T. Heath. Computing a sparse basis for the null space. SIAM Journal on Algebraic and Discrete Methods, 8(3):446–459, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Griggs:1986:SAM

[GHW86] Jerrold R. Griggs, Philip J. Hanlon, and Michael S. Waterman. Sequence alignments with matched sections. SIAM Journal on Algebraic and Discrete Methods, 7(4):604–608, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Gilbert:1980:RCL

[Gil80a] E. N. Gilbert. Random colorings of a lattice of squares in the plane. SIAM Journal on Algebraic and Discrete Methods, 1 (2):152–159, June 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Gilbert:1980:NNC

[Gil80b] John R. Gilbert. A note on the NP-completeness of vertex elimination on directed graphs. SIAM Journal on Algebraic and Discrete Methods, 1(3):292–294, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Giles:1981:APP

[Gil81] Rick Giles. Adjacency on the postman polyhedron. SIAM Journal on Algebraic and Discrete Methods, 2(2):172–175, June 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Garey:1983:CNN

[GJ83] M. R. Garey and D. S. Johnson. Crossing number is NPcomplete. SIAM Journal on Algebraic and Discrete Methods, 4(3):312–316, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Goulden:1986:LGS

[GJ86] I. P. Goulden and D. M. Jackson. Labelled graphs with small vertex degrees and P - recursiveness. SIAM Journal on Algebraic and Discrete Methods, 7(1):60–66, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Garey:1980:CCC

[GJMP80] M. R. Garey, D. S. Johnson, G. L. Miller, and C. H. Papadimitriou. The complexity of coloring circular arcs and chords. SIAM Journal on Algebraic and Discrete Methods, 1(2):216–227, June 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Goulden:1983:HSS

[GJR83] I. P. Goulden, D. M. Jackson, and J. W. Reilly. The Hammond series of a symmetric function and its application to P - recursiveness. SIAM Journal on Algebraic and Discrete Methods, 4(2):179–193, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Garey:1983:SOF

[GJTY83] M. R. Garey, D. S. Johnson, R. E. Tarjan, and M. Yannakakis. Scheduling opposing forests. SIAM Journal on Algebraic and Discrete Methods, 4 (1):72–93, March 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Graham:1980:SD

[GLL80] R. L. Graham, S.-Y. R. Li, and W. C. W. Li. On the structure of t-designs. SIAM Journal on Algebraic and Discrete Methods, 1(1):8–14, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Greenberg:1981:GTM

[GLM81] Harvey J. Greenberg, J. Richard Lundgren, and John S. Maybee. Graph theoretic methods for the qualitative analysis of rectangular matrices. SIAM Journal on Algebraic and Discrete Methods, 2(3):227–239, September 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Greenberg:1983:RMS

[GLM83] Harvey J. Greenberg, J. Richard Lundgren, and John S. Maybee. Rectangular matrices and signed graphs. SIAM Journal on Algebraic and Discrete Methods, 4(1):50–61, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Greenberg:1984:ISG

[GLM84] Harvey J. Greenberg, J. Richard Lundgren, and John S. Maybee. Inverting signed graphs. SIAM Journal on Algebraic and Discrete Methods, 5(2):216–223, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Golub:1986:UFG

[GM86] Gene H. Golub and Carl D.

Meyer, Jr. Using the QR factorization and group inversion to compute, differentiate, and estimate the sensitivity of stationary probabilities for Markov chains. SIAM Journal on Algebraic and Discrete Methods, 7(2):273–281, April 1986. CODEN SJAMDU.

ISSN 0196-5212 (print), 2168-345X (electronic).

Gibbons:1987:ESF

[GM87a] P. B. Gibbons and E. Mendelsohn. The existence of a subsquare free Latin square of side 12. SIAM Journal on Algebraic and Discrete Methods, 8(1):93– 99, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Green:1987:FVF

[GM87b] W. L. Green and T. D. Morley. Four variational formulations of the contraharmonic mean of operators. SIAM Journal on Algebraic and Discrete Methods, 8(4):670–673, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Golomb:1980:OSI

[Gol80] Solomon W. Golomb. Obtaining specified irreducible polynomials over finite fields. SIAM Journal on Algebraic and Discrete Methods, 1(4):411–418, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Goles:1982:FPB

[Gol82] Eric Goles. Fixed point behavior of threshold functions on a finite set. SIAM Journal on Algebraic and Discrete Methods, 3(4):529– 531, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Goldman:1985:PUM

[Gol85a] Ronald N. Goldman. Polya’s urn model and computer aided geometric design. SIAM Journal on Algebraic and Discrete Methods, 6(1):1–28, January 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

GolesCh:1985:DBN

[Gol85b] E. Goles Ch. Dynamical behaviour of neural networks. SIAM Journal on Algebraic and Discrete Methods, 6(4):749–754, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Gierz:1983:MSO

[GP83] Gerhard Gierz and Werner Poguntke. Minimizing setups for ordered sets: a linear algebraic approach. SIAM Journal on Algebraic and Discrete Methods, 4 (1):132–144, March 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Guan:1985:EOC

[GP85] Mei Gu Guan and William Pulleyblank. Eulerian orientations and circulations. SIAM Journal on Algebraic and Discrete Methods, 6(4):657–664, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Granovsky:1981:OFC

[Gra81] B. L. Granovsky. Optimal formulae of the conditional Monte Carlo. SIAM Journal on Algebraic and Discrete Methods, 2(3):289–294, September 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Gilbert:1984:STC

[GRE84] John R. Gilbert, Donald J. Rose, and Anders Edenbrandt. A separator theorem for chordal graphs. SIAM Journal on Algebraic and Discrete Methods, 5(3):306–313, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Grunbaum:1981:ETM

[Grü81] F. Alberto Grünbaum. Eigenvectors of a Toeplitz matrix: discrete version of the prolate spheroidal wave functions. SIAM Journal on Algebraic and Discrete Methods, 2(2):136–141, 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Grunbaum:1986:PLD

[Grü86] F. Alberto Grünbaum. A property of the Legendre differential equation and its discretization. SIAM Journal on Algebraic and Discrete Methods, 7(3):358– 361, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Graham:1980:ABH

[GS80] R. L. Graham and N. J. A. Sloane. On additive bases and harmonious graphs. SIAM Journal on Algebraic and Discrete Methods, 1(4):382–404, 1980.

CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Garfinkel:1984:HCS

[GS84] R. S. Garfinkel and P. S. Sundararaghavan. Hamiltonian cycles in striped graphs: the twostripe problem. SIAM Journal on Algebraic and Discrete Methods, 5(4):463–466, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Gupta:1986:SET

[Gup86] Murli M. Gupta. A spectrum enveloping technique for iterative solution of central difference approximations of convectiondiffusion equations. SIAM Journal on Algebraic and Discrete Methods, 7(4):513–526, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Griggs:1980:EVI

[GW80] Jerrold R. Griggs and Douglas B.

West. Extremal values of the interval number of a graph. SIAM Journal on Algebraic and Discrete Methods, 1(1):1–7, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Goldberg:1985:BCC

[GW85] Charles H. Goldberg and Douglas B. West. Bisection of circle colorings. SIAM Journal on Algebraic and Discrete Methods, 6(1):93–106, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Gyori:1986:SPR

[Gyö86] Ervin Györi. A short proof of the rectilinear art gallery theorem. SIAM Journal on Algebraic and Discrete Methods, 7(3):452– 454, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Graham:1980:SMP

[GYY80] R. L. Graham, A. C. Yao, and F. F. Yao. Some monotonicity properties of partial orders. SIAM Journal on Algebraic and Discrete Methods, 1(3):251– 258, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hadjidimos:1985:OCI

[Had85] A. Hadjidimos. On the optimization of the classical iterative schemes for the solution of complex singular linear systems. SIAM Journal on Algebraic and Discrete Methods, 6(4):555–566, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hall:1985:NSN

[Hal85] Charles A. Hall. Numerical solution of Navier–Stokes problems by the dual variable method. SIAM Journal on Algebraic and Discrete Methods, 6(2):220–236, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Harrod:1986:DTI

[Har86] W. J. Harrod. LU-decompositions of tridiagonal irreducible Hmatrices. SIAM Journal on Algebraic and Discrete Methods, 7(2):180–187, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hashimoto:1985:DFM

[Has85] Hiroshi Hashimoto. Decomposition of fuzzy matrices. SIAM Journal on Algebraic and Discrete Methods, 6(1):32–38, January 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Haviv:1986:ASDb

[Hav86] Moshe Haviv. An approximation to the stationary distribution of a nearly completely decomposable Markov chain and its error analysis. SIAM Journal on Algebraic and Discrete Methods, 7(4):589–593, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Heath:1987:EOG

[Hea87] Lenwood S. Heath. Embedding outerplanar graphs in small books. SIAM Journal on Algebraic and Discrete Methods, 8(2):198–218, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Herden:1984:AAC

[Her84a] Gerhard Herden. The approximation of arbitrary clustering functions by clustering functions which satisfy overlapping criteria. SIAM Journal on Algebraic and Discrete Methods, 5(4):497– 507, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Herden:1984:SAC

[Her84b] Gerhard Herden. Some aspects of clustering functions. SIAM Journal on Algebraic and Discrete Methods, 5(1):101–116, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Heyman:1987:FCD

[Hey87] Daniel P. Heyman. Further comparisons of direct methods for computing stationary distributions of Markov chains. SIAM Journal on Algebraic and Discrete Methods, 8(2):226–232, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Homer:1985:DPS

[HG85] Steven Homer and Jerry Goldman. Doubly-periodic sequences and two-dimensional recurrences. SIAM Journal on Algebraic and Discrete Methods, 6(3):360–370, July 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Haggkvist:1982:SMR

[HH82] R. Häggkvist and P. Hell. Sorting and merging in rounds. SIAM Journal on Algebraic and Discrete Methods, 3(4):465– 473, December 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hong:1986:RMT

[HH86] Yoo Pyo Hong and Roger A.

Horn. On the reduction of a matrix to triangular or diagonal form by consimilarity. SIAM Journal on Algebraic and Discrete Methods, 7(1):80–88, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hammer:1982:VBA

[HHS82] P. L. Hammer, P. Hansen, and B. Simeone. Vertices belonging to all or to no maximum stable sets of a graph. SIAM Journal on Algebraic and Discrete Methods, 3(4):511–522, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hu:1981:BPG

[HHW81] M. C. Hu, F. K. Hwang, and Ju Kwei Wang. A boundary problem for group testing. SIAM Journal on Algebraic and Discrete Methods, 2(2):81–87, June 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hilliard:1987:CEH

[Hil87] Lorenzo O. Hilliard. The case of equality in Hopf’s inequality. SIAM Journal on Algebraic and Discrete Methods, 8(4):691– 709, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hammer:1981:TS

[HIS81] P. L. Hammer, T. Ibaraki, and B. Simeone. Threshold sequences. SIAM Journal on Algebraic and Discrete Methods, 2 (1):39–49, March 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hopcroft:1983:HCG

[HK83] J. E. Hopcroft and M. S. Krishnamoorthy. On the harmonious coloring of graphs. SIAM Journal on Algebraic and Discrete Methods, 4(3):306–311, September 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic). Work related to minimal perfect hash functions.

Hell:1986:PCB

[HK86] P. Hell and D. G. Kirkpatrick. Packings by complete bipartite graphs. SIAM Journal on Algebraic and Discrete Methods, 7(2):199–209, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hoffman:1985:TBG

[HKS85] A. J. Hoffman, A. W. J. Kolen, and M. Sakarovitch. Totallybalanced and greedy matrices. SIAM Journal on Algebraic and Discrete Methods, 6(4):721–730, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Han:1984:CPD

[HM84] S.-P. Han and O. L. Mangasarian. Characterization of positive definite and semidefinite matrices via quadratic programming duality. SIAM Journal on Algebraic and Discrete Methods, 5 (1):26–32, March 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hammer:1985:BG

[HM85] P. L. Hammer and N. V. R.

Mahadev. Bithreshold graphs. SIAM Journal on Algebraic and Discrete Methods, 6(3):497–506, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hershkowitz:1987:MSS

[HMS87] Daniel Hershkowitz, Volker Mehrmann, and Hans Schneider. Matrices with sign symmetric diagonal shifts or scalar shifts. SIAM Journal on Algebraic and Discrete Methods, 8(1):108–122, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hochbaum:1983:FSS

[Hoc83] Dorit S. Hochbaum. On the fractional solution to the set covering problem. SIAM Journal on Algebraic and Discrete Methods, 4(2):221–222, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Helman:1985:CMD

[HR85] Paul Helman and Arnon Rosenthal. A comprehensive model of dynamic programming. SIAM Journal on Algebraic and Discrete Methods, 6(2):319–334, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Haviv:1986:ASDa

[HR86] Moshe Haviv and Y. Ritov. An approximation to the stationary distribution of a nearly completely decomposable Markov chain and its error bound. SIAM Journal on Algebraic and Discrete Methods, 7(4):583–588, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Haddad:1987:GFT

[HRS87] Ramsey W. Haddad, Shaibal Roy, and Alejandro A. Schäffer. On gossiping with faulty telephone lines. SIAM Journal on Algebraic and Discrete Methods, 8(3):439–445, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hulme:1981:MLD

[HS81] B. L. Hulme and P. J. Slater. Minimean location of different facilities on a line network. SIAM Journal on Algebraic and Discrete Methods, 2(4):411–415, December 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hochbaum:1985:APC

[HS85] Dorit S. Hochbaum and David B. Shmoys. An O(|V |^{2}) algorithm for the planar 3-cut problem. SIAM Journal on Algebraic and Discrete Methods, 6(4):707– 712, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hochbaum:1986:PPY

[HS86] Dorit S. Hochbaum and David B.

Shmoys. A packing problem you can almost solve by sitting on your suitcase. SIAM Journal on Algebraic and Discrete Methods, 7(2):247–257, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hwang:1981:HGH

[HSD81] F. K. Hwang, Tien Tai Song, and Ding Zhu Du. Hypergeometric and generalized hypergeometric group testing. SIAM Journal on Algebraic and Discrete Methods, 2(4):426–428, December 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hwang:1985:OSP

[HSY85] F. K. Hwang, J. Sun, and E. Y.

Yao. Optimal set partitioning. SIAM Journal on Algebraic and Discrete Methods, 6 (1):163–170, January 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hassin:1986:EAO

[HT86] R. Hassin and A. Tamir. Efficient algorithms for optimization and selection on seriesparallel graphs. SIAM Journal on Algebraic and Discrete Methods, 7(3):379–389, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hill:1985:MAA

[HU85] Richard D. Hill and E. Eugene Underwood. On the matrix adjoint (adjugate). SIAM Journal on Algebraic and Discrete Methods, 6(4):731–737, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hulme:1984:BMO

[Hul84] Bernie L. Hulme. Boolean methods of optimization over independence systems. SIAM Journal on Algebraic and Discrete Methods, 5(2):255–262, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Harary:1982:PSG

[HVW82] Frank Harary, Andrew Vince, and Dale Worley. A pointsymmetric graph that is nowhere reversible. SIAM Journal on Algebraic and Discrete Methods, 3(3):285–287, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hwang:1984:TVG

[Hwa84] F. K. Hwang. Three versions of a group testing game. SIAM Journal on Algebraic and Discrete Methods, 5(2):145–153, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Hwang:1986:SRC

[Hwa86] F. K. Hwang. Simplified reliabilities for consecutive-k-out-of-n systems. SIAM Journal on Algebraic and Discrete Methods, 7(2):258–264, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Inukai:1981:WCM

[IW81] Thomas Inukai and Louis Weinberg. Whitney connectivity of matroids. SIAM Journal on Algebraic and Discrete Methods, 2 (2):108–120, June 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Johnson:1987:FPG

[JM87] Ellis L. Johnson and Sebastiano Mosterts. On four problems in graph theory. SIAM Journal on Algebraic and Discrete Methods, 8(2):163–185, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Johnson:1985:CNS

[Joh85] Carolyn R. Johnson. Convergence of a nonlinear sharpening transformation for digital images. SIAM Journal on Algebraic and Discrete Methods, 6 (3):462–465, July 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Jones:1982:COI

[Jon82] Philip C. Jones. Computing an optimal invariant capital stock. SIAM Journal on Algebraic and Discrete Methods, 3(2):145–150, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Johnson:1983:CMT

[JOv83] Charles R. Johnson, D. D. Olesky, and P. van den Driessche. A class of M-matrices with tree graphs. SIAM Journal on Algebraic and Discrete Methods, 4(4):476–480, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Johnson:1985:CSH

[JR85] Charles R. Johnson and Leiba Rodman. Convex sets of Hermitian matrices with constant inertia. SIAM Journal on Algebraic and Discrete Methods, 6(3):351– 359, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Jain:1981:NMM

[JS81] S. K. Jain and L. E. Snyder. Nonnegative λ-monotone matrices. SIAM Journal on Algebraic and Discrete Methods, 2 (1):66–76, March 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Johnson:1986:MAR

[JS86] Charles R. Johnson and Helene M. Shapiro. Mathematical aspects of the relative gain array (A ∘ A−T). SIAM Journal on Algebraic and Discrete Methods, 7(4):627–644, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kantor:1982:STPb

[Kan82a] W. M. Kantor. Spreads, translation planes and Kerdock sets. II. SIAM Journal on Algebraic and Discrete Methods, 3(3):308– 318, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kantor:1982:STPa

[Kan82b] William M. Kantor. Spreads, translation planes and Kerdock sets. I. SIAM Journal on Algebraic and Discrete Methods, 3(2):151–165, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Klawe:1982:ITH

[KCP82] M. M. Klawe, D. G. Corneil, and A. Proskurowski. Isomorphism testing in hookup classes. SIAM Journal on Algebraic and Discrete Methods, 3(2):260–274, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kirschenhofer:1983:SOA

[Kir83] Peter Kirschenhofer. Subblock occurrences in the q-ary representation of n. SIAM Journal on Algebraic and Discrete Methods, 4(2):231–236, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kahn:1983:TGR

[KKK83] J. Kahn, M. Klawe, and D. Kleitman. Traditional galleries require fewer watchmen. SIAM Journal on Algebraic and Discrete Methods, 4(2):194–206, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kirkpatrick:1985:SGC

[KKP85] David G. Kirkpatrick, Maria Klawe, and Nicholas Pippenger. Some graph-colouring theorems with applications to generalized connection networks. SIAM Journal on Algebraic and Discrete Methods, 6(4):576–582, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kaltofen:1987:FPC

[KKS87] Erich Kaltofen, M. S. Krishnamoorthy, and B. David Saunders. Fast parallel computation of Hermite and Smith forms of polynomial matrices. SIAM Journal on Algebraic and Discrete Methods, 8(4):683–690, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Korte:1984:GLO

[KL84] Bernhard Korte and László Lovász. Greedoids and linear objective functions. SIAM Journal on Algebraic and Discrete Methods, 5(2):229–238, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Koury:1984:IMC

[KMS84] J. R. Koury, D. F. McAllister, and W. J. Stewart. Iterative methods for computing stationary distributions of nearly completely decomposable Markov chains. SIAM Journal on Algebraic and Discrete Methods, 5(2):164–186, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Korner:1986:FKB

[Kör86] János Körner. Fredman-Komlós bounds and information theory. SIAM Journal on Algebraic and Discrete Methods, 7(4):560– 570, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kincaid:1986:VCS

[KOY86] David R. Kincaid, Thomas C.

Oppe, and David M. Young. Vector computations for sparse linear systems. SIAM Journal on Algebraic and Discrete Methods, 7(1):99–112, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Krishna:1985:EBS

[Kri85] Lala B. Krishna. Error bounds for the SSOR semi-iterative method. SIAM Journal on Algebraic and Discrete Methods, 6(2):245–250, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kleitman:1981:SOR

[KS81] D. J. Kleitman and Michael E.

Saks. Set orderings requiring costliest alphabetic binary trees. SIAM Journal on Algebraic and Discrete Methods, 2(2):142–146, June 1981. CODEN SJAMDU.

ISSN 0196-5212 (print), 2168-345X (electronic).

Kano:1985:RVP

[KS85] Mikio Kano and Akio Sakamoto. Ranking the vertices of a paired comparison digraph. SIAM Journal on Algebraic and Discrete Methods, 6(1):79–92, January 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kilby:1987:CRPa

[KS87a] Karen E. Kilby and N. J. A. Sloane. On the covering radius problem for codes. I. Bounds on normalized covering radius. SIAM Journal on Algebraic and Discrete Methods, 8(4):604–618, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kilby:1987:CRPb

[KS87b] Karen E. Kilby and N. J. A. Sloane. On the covering radius problem for codes. II. Codes of low dimension; normal and abnormal codes. SIAM Journal on Algebraic and Discrete Methods, 8(4):619–627, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kundu:1980:RPG

[KSSS80] Sukhamay Kundu, E. Sampathkumar, James Shearer, and Dean Sturtevant. Reconstruction of a pair of graphs from their concatenations. SIAM Journal on Algebraic and Discrete Methods, 1(2):228–231, June 1980.

CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kelly:1981:ENV

[KT81] D. G. Kelly and J. W. Tolle.

Expected number of vertices of a random convex polyhedron. SIAM Journal on Algebraic and Discrete Methods, 2(4):441–451, December 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kerov:1986:CIS

[KV86] Sergei V. Kerov and Anatol M.

Vershik. The characters of the infinite symmetric group and probability properties of the Robinson-Schensted-Knuth algorithm. SIAM Journal on Algebraic and Discrete Methods, 7(1):116–124, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kanetkar:1980:CMD

[KW80a] Sharad V. Kanetkar and Meghanad D.

Wagh. On construction of matrices with distinct submatrices. SIAM Journal on Algebraic and Discrete Methods, 1(1):107– 113, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kleitman:1980:ELS

[KW80b] Daniel J. Kleitman and Kenneth J. Winston. Ensembles and largest sojourns of random walks. SIAM Journal on Algebraic and Discrete Methods, 1(1):15–20, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Lamond:1987:EFG

[Lam87] Bernard F. Lamond. An efficient factorization for the group inverse. SIAM Journal on Algebraic and Discrete Methods, 8(4):797–808, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Langston:1984:PHC

[Lan84] Michael A. Langston. Performance of heuristics for a computer resource allocation problem. SIAM Journal on Algebraic and Discrete Methods, 5 (2):154–161, June 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Leibowitz:1982:ICG

[LAP82] R. Leibowitz, S. F. Assmann, and G. W. Peck. The interval count of a graph. SIAM Journal on Algebraic and Discrete Methods, 3(4):485–494, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Latouche:1982:PTS

[Lat82]  Guy Latouche.  A phase-

type semi-Markov point process. SIAM Journal on Algebraic and Discrete Methods, 3(1):77–90, March 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Liu:1984:OFS

[LC84] C. J. Liu and Yutze Chow.

On operator and formal sum methods for graph enumeration problems. SIAM Journal on Algebraic and Discrete Methods, 5(3):384–406, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Lengauer:1982:ULB

[Len82] Thomas Lengauer. Upper and lower bounds on the complexity of the min-cut linear arrangement problem on trees. SIAM Journal on Algebraic and Discrete Methods, 3(1):99–113, March 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Lewin:1980:TNJ

[Lew80] Mordechai Lewin. Totally nonnegative, M-, and Jacobi matrices. SIAM Journal on Algebraic and Discrete Methods, 1(4):419– 421, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Lih:1982:MFP

[Lih82]  Ko Wei Lih.  Majorization

of finite partially ordered sets. SIAM Journal on Algebraic and Discrete Methods, 3(4):495–503, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Linial:1986:HEP

[Lin86] Nathan Linial. Hard enumeration problems in geometry and combinatorics. SIAM Journal on Algebraic and Discrete Methods, 7(2):331–335, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Lingas:1987:NHM

[Lin87] Andrzej Lingas. A new heuristic for minimum weight triangulation. SIAM Journal on Algebraic and Discrete Methods, 8(4):646– 658, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Liu:1987:AGT

[Liu87] Joseph W. H. Liu. An application of generalized tree pebbling to sparse matrix factorization. SIAM Journal on Algebraic and Discrete Methods, 8(3):375– 395, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Lewandowski:1987:STS

[LL87] J. L. Lewandowski and C. L. Liu. SS/TDMA satellite communications with k-permutation switching modes. SIAM Journal on Algebraic and Discrete Methods, 8(4):519–534, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Leighton:1984:LSE

[LLM84] Frank Thomson Leighton, Margaret Lepley, and Gary L. Miller. Layouts for the shuffle-exchange graph based on the complex plane diagram. SIAM Journal on Algebraic and Discrete Methods, 5(2):202–215, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Lucas:1982:FST

[LM82] William F. Lucas and Kai Michaelis. Finite solution theory for coalitional games. SIAM Journal on Algebraic and Discrete Methods, 3(4):551–565, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Latouche:1980:EAS

[LN80] Guy Latouche and Marcel F.

Neuts. Efficient algorithmic solutions to exponential tandem queues with blocking. SIAM Journal on Algebraic and Discrete Methods, 1(1):93–106, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Laskar:1984:ACT

[LPHH84] Renu Laskar, John Pfaff, S. M.

Hedetniemi, and S. T. Hedetniemi. On the algorithmic complexity of total domination. SIAM Journal on Algebraic and Discrete Methods, 5(3):420–425, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Linial:1981:IMS

[LR81] Nathan Linial and Bruce L.

Rothschild. Incidence matrices of subsets—a rank formula. SIAM Journal on Algebraic and Discrete Methods, 2(3):333–340, September 1981.

CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Lancaster:1983:ME

[LR83] P. Lancaster and P. Rózsa. On the matrix equation AX + X^{∗}A^{∗} = C. SIAM Journal on Algebraic and Discrete Methods, 4(4):432–436, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Letac:1980:RWC

[LT80] Gérard Letac and Lajos Takács. Random walks on a 600-cell. SIAM Journal on Algebraic and Discrete Methods, 1(1):114–120, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Lovasz:1982:GRP

[LY82] L. Lovász and Y. Yemini. On generic rigidity in the plane. SIAM Journal on Algebraic and Discrete Methods, 3(1):91–98, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Mack:1981:LPN

[Mac81] Y. P. Mack. Local properties of k-NN regression estimates. SIAM Journal on Algebraic and Discrete Methods, 2(3):311–323, September 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Mattheij:1984:SBD

[Mat84] R. M. M. Mattheij. Stability of block LU-decompositions of matrices arising from BVP. SIAM Journal on Algebraic and Discrete Methods, 5(3):314–331, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Matthews:1987:MRR

[Mat87] Peter Matthews. Mixing rates for a random walk on the cube. SIAM Journal on Algebraic and Discrete Methods, 8(4):746–752, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Mitchison:1986:ONA

[MD86] Graeme Mitchison and Richard Durbin. Optimal numberings of an N × N array. SIAM Journal on Algebraic and Discrete Methods, 7(4):571–582, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Mehravari:1986:GBB

[Meh86]  Nader Mehravari.  General-

ized binary binomial group testing. SIAM Journal on Algebraic and Discrete Methods, 7(1):159– 166, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Merris:1986:SIP

[Mer86] Russell Merris. The second immanantal polynomial and the centroid of a graph. SIAM Journal on Algebraic and Discrete Methods, 7(3):484–503, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Meyer:1980:CFM

[Mey80] Carl D. Meyer, Jr. The condition of a finite Markov chain and perturbation bounds for the limiting probabilities. SIAM Journal on Algebraic and Discrete Methods, 1(3):273–283, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Minamide:1985:EMI

[Min85a] Nariyasu Minamide. An extension of the matrix inversion lemma. SIAM Journal on Algebraic and Discrete Methods, 6(3):371–377, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Minamide:1985:RBA

[Min85b] Nariyasu Minamide. Recursive best approximate solution algorithms. SIAM Journal on Algebraic and Discrete Methods, 6 (3):378–383, July 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Murota:1987:CCF

[MIN87] Kazuo Murota, Masao Iri, and Masataka Nakamura. Combinatorial canonical form of layered mixed matrices and its application to block-triangularization of systems of linear/nonlinear equations. SIAM Journal on Algebraic and Discrete Methods, 8(1):123–149, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Meir:1980:APO

[MMP80a] A. Meir, J. W. Moon, and J. R.

Pounder. Acknowledgment of priority: “On the order of random channel networks”. SIAM Journal on Algebraic and Discrete Methods, 1(4):462, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Meir:1980:ORC

[MMP80b] A. Meir, J. W. Moon, and J. R.

Pounder. On the order of random channel networks. SIAM Journal on Algebraic and Discrete Methods, 1(1):25–33, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Magnus:1980:EMS

[MN80] Jan R. Magnus and H. Neudecker.

The elimination matrix: some lemmas and applications. SIAM Journal on Algebraic and Discrete Methods, 1(4):422–449, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Moses:1984:ERG

[MN84] Yoram Moses and Amos Noy.

On the encoding of relations by graphs. SIAM Journal on Algebraic and Discrete Methods, 5(4):569–578, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Mollard:1986:GPF

[Mol86] M. Mollard. A generalized parity function and its use in the construction of perfect codes. SIAM Journal on Algebraic and Discrete Methods, 7(1):113–115, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Monien:1986:BMP

[Mon86] Burkhard Monien. The bandwidth minimization problem for caterpillars with hair length 3 is NP-complete. SIAM Journal on Algebraic and Discrete Methods, 7(4):505–512, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Moon:1983:LNA

[Moo83] J. W. Moon. On level numbers of t-ary trees. SIAM Journal on Algebraic and Discrete Methods, 4(1):8–13, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Morrison:1986:WAR

[Mor86] J. A. Morrison. Weighted averages of Radon transforms on Z_{2}^{k}. SIAM Journal on Algebraic and Discrete Methods, 7(3):404– 413, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Mahmoud:1984:MPS

[MP84a] Hosam Mahmoud and Boris Pittel. On the most probable shape of a search tree grown from a random permutation.

SIAM Journal on Algebraic and Discrete Methods, 5(1):69–81, March 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Moody:1984:CEF

[MP84b] R. V. Moody and J. Patera.

Characters of elements of finite order in Lie groups. SIAM Journal on Algebraic and Discrete Methods, 5(3):359–383, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Makedon:1985:TB

[MPS85] F. S. Makedon, C. H. Papadimitriou, and I. H. Sudborough. Topological bandwidth. SIAM Journal on Algebraic and Discrete Methods, 6(3):418–444, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Mallows:1981:WES

[MS81] C. L. Mallows and N. J. A.

Sloane. Weight enumerators of self-orthogonal codes over GF(3). SIAM Journal on Algebraic and Discrete Methods, 2(4):452–460, December 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Mangasarian:1986:VCN

[MS86] O. L. Mangasarian and T. H.

Shiau. A variable-complexity norm maximization problem. SIAM Journal on Algebraic and Discrete Methods, 7(3):455–461, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Megiddo:1983:FLD

[MT83] Nimrod Megiddo and Arie Tamir. Finding least-distances lines. SIAM Journal on Algebraic and Discrete Methods, 4 (2):207–211, June 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

McKay:1981:CE

[MW81] James H. McKay and Stuart Sui-Sheng Wang. The v×v (0, 1, −1)- circulant equation AAT = vI − J. SIAM Journal on Algebraic and Discrete Methods, 2(3):266– 274, September 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Merris:1983:TG

[MW83] Russell Merris and William Watkins. Tensors and graphs. SIAM Journal on Algebraic and Discrete Methods, 4(4):534–547, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Maybee:1987:FTI

[MW87] John S. Maybee and Gerry M. Wiener. L-functions and their inverses. SIAM Journal on Algebraic and Discrete Methods, 8(1):67–76, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Megiddo:1983:MCL

[MZH83] Nimrod Megiddo, Eitan Zemel, and S. Louis Hakimi. The maximum coverage location problem. SIAM Journal on Algebraic and Discrete Methods, 4(2):253–261, June 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Neumaier:1983:LST

[Neu83] A. Neumaier. Lattices of simplex type. SIAM Journal on Algebraic and Discrete Methods, 4(2):145–160, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Neumann:1986:NDC

[Neu86] Michael Neumann. Neighborhoods of dominant convergence for the SSOR method. SIAM Journal on Algebraic and Discrete Methods, 7(4):551–559, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Newsam:1983:ESJ

[NR83] Garry N. Newsam and John D.

Ramsdell. Estimation of sparse Jacobian matrices. SIAM Journal on Algebraic and Discrete Methods, 4(3):404–418, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Nain:1981:FRM

[NS81] R. B. Nain and Kanwar Sen.

Fluctuation results for Markovdependent trials. SIAM Journal on Algebraic and Discrete Methods, 2(4):341–346, December 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Orlin:1981:ACP

[OBB81] James B. Orlin, Maurizio A. Bonuccelli, and Daniel P. Bovet. An O(n^{2}) algorithm for coloring proper circular arc graphs. SIAM Journal on Algebraic and Discrete Methods, 2(2):88–93, June 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Dokovic:1987:QCI

[– Dok87] Dragomir Ž. – Doković. Quadratic cones invariant under some linear operators. SIAM Journal on Algebraic and Discrete Methods, 8(2):186–191, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Opsut:1982:CCN

[Ops82] Robert J. Opsut. On the computation of the competition number of a graph. SIAM Journal on Algebraic and Discrete Methods, 3(4):420–428, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

OLeary:1985:MMP

[OW85] Dianne P. O’Leary and R. E. White. Multisplittings of matrices and parallel solution of linear systems. SIAM Journal on Algebraic and Discrete Methods, 6(4):630–640, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Owen:1986:VGR

[Owe86] Guillermo Owen. Values of graph-restricted games. SIAM Journal on Algebraic and Discrete Methods, 7(2):210–220, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Peck:1981:HTS

[Pec81] G. W. Peck. A Helly theorem for sets. SIAM Journal on Algebraic and Discrete Methods, 2(3):306–308, September 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Perlstadt:1981:SSF

[Per81]  Marci Perlstadt.  Sampling

schemes for Fourier transform reconstruction. SIAM Journal on Algebraic and Discrete Methods, 2(2):176–191, June 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Perlstadt:1983:COP

[Per83] Marci Perlstadt. Chopped orthogonal polynomial expansions— some discrete cases. SIAM Journal on Algebraic and Discrete Methods, 4(1):94–100, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Perline:1987:DTB

[Per87] Ronald Keith Perline. Discrete time-band limiting operators and commuting tridiagonal matrices. SIAM Journal on Algebraic and Discrete Methods, 8(2):192–195, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Phelps:1983:CCP

[Phe83] K. T. Phelps. A combinatorial construction of perfect codes. SIAM Journal on Algebraic and Discrete Methods, 4(3):398–403, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Phelps:1984:GPC

[Phe84] K. T. Phelps. A general product construction for error correcting codes. SIAM Journal on Algebraic and Discrete Methods, 5(2):224–228, June 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Philippe:1987:AIN

[Phi87] Bernard Philippe. An algorithm to improve nearly orthonormal sets of vectors on a vector processor. SIAM Journal on Algebraic and Discrete Methods, 8(3):396– 403, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Pippenger:1980:NLB

[Pip80] Nicholas Pippenger. A new lower bound for the number of switches in rearrangeable networks. SIAM Journal on Algebraic and Discrete Methods, 1 (2):164–167, June 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Plemmons:1986:PBI

[Ple86] Robert J. Plemmons. A parallel block iterative scheme applied to computations in structural analysis. SIAM Journal on Algebraic and Discrete Methods, 7(3):337–347, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Plumstead:1985:BCC

[PP85] Barton R. Plumstead and Joan B. Plumstead. Bounds for cube coloring. SIAM Journal on Algebraic and Discrete Methods, 6(1):73–78, January 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Peschon:1982:NME

[PR82] J. Peschon and D. Ross. New methods for evaluating distribution automation and control (DAC) systems benefits. SIAM Journal on Algebraic and Discrete Methods, 3(4):439– 452, December 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Peskin:1985:MCM

[PR85] Barbara R. Peskin and David R.

Richman. A method to compute minimal polynomials. SIAM Journal on Algebraic and Discrete Methods, 6(2):292–299, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Proctor:1982:RPS

[Pro82a] Robert A. Proctor. Representations of sl(2, C) on posets and the Sperner property. SIAM Journal on Algebraic and Discrete Methods, 3(2):275–280, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Prodinger:1982:GSD

[Pro82b] Helmut Prodinger. Generalizing the sum of digits function. SIAM Journal on Algebraic and Discrete Methods, 3 (1):35–42, March 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Provan:1983:DLS

[Pro83] J. Scott Provan. Determinacy in linear systems and networks. SIAM Journal on Algebraic and Discrete Methods, 4(2):262–278, June 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Proctor:1985:IDB

[Pro85] P. E. Proctor. Implementation of a double-basis simplex method for the general linear programming problem. SIAM Journal on Algebraic and Discrete Methods, 6(4):567–575, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Provan:1987:SCC

[Pro87] J. Scott Provan. Substitutes and complements in constrained linear models. SIAM Journal on Algebraic and Discrete Methods, 8(4):585–603, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Perl:1983:EOM

[PS83] Yehoshua Perl and Yossi Shiloach.

Efficient optimization of monotonic functions on trees. SIAM Journal on Algebraic and Discrete Methods, 4(4):512–516, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Proskurowski:1986:EVE

[PS86] Andrzej Proskurowski and Maciej M. Sysło. Efficient vertexand edge-coloring of outerplanar graphs. SIAM Journal on Algebraic and Discrete Methods, 7(1):131–136, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Pippenger:1982:RNL

[PY82] Nicholas Pippenger and Andrew C.-C. C. Yao. Rearrangeable networks with limited depth. SIAM Journal on Algebraic and Discrete Methods, 3(4):411–417, December 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Rabie:1985:SGN

[Rab85] Mohamed A. Rabie. A simple game with no symmetric solution. SIAM Journal on Algebraic and Discrete Methods, 6 (1):29–31, January 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Randic:1985:CEC

[Ran85] Milan Randić. On the characteristic equations of the characteristic polynomial. SIAM Journal on Algebraic and Discrete Methods, 6(1):145–162, January 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Redheffer:1985:VMIa

[Red85a] Ray Redheffer. Volterra multipliers. I. SIAM Journal on Algebraic and Discrete Methods, 6(4):592–611, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Redheffer:1985:VMIb

[Red85b] Ray Redheffer. Volterra multipliers. II. SIAM Journal on Algebraic and Discrete Methods, 6(4):612–623, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Redheffer:1986:EVM

[Red86]  Ray Redheffer.  Erratum:

“Volterra multipliers. II”. SIAM Journal on Algebraic and Discrete Methods, 7(2):336, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Ross:1982:EPN

[Ros82] Jeffrey A. Ross. On the exponent of a primitive, nearly reducible matrix. II. SIAM Journal on Algebraic and Discrete Methods, 3(3):395–410, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Rose:1984:CRS

[Ros84] Donald J. Rose. Convergent regular splittings for singular Mmatrices. SIAM Journal on Algebraic and Discrete Methods, 5(1):133–144, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Rothblum:1981:SGA

[Rot81] Uriel G. Rothblum. Sensitive growth analysis of multiplicative systems. I. The dynamic approach. SIAM Journal on Algebraic and Discrete Methods, 2 (1):25–34, March 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Rothblum:1984:ESO

[Rot84] Uriel G. Rothblum. Explicit solutions to optimization problems on the intersections of the unit ball of the l_{1} and l_{∞} norms with a hyperplane. SIAM Journal on Algebraic and Discrete Methods, 5(4):619–632, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Rabinowitz:1985:AAC

[RP85] Joshua H. Rabinowitz and Viera Krňanová Proulx. An asymptotic approach to the channel assignment problem. SIAM Journal on Algebraic and Discrete Methods, 6(3):507–518, July 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Rosenblatt:1982:SHS

[RS82] Joseph Rosenblatt and Paul D. Seymour. The structure of homometric sets. SIAM Journal on Algebraic and Discrete Methods, 3(3):343–350, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Robertson:1985:DPS

[RS85] Neil Robertson and P. D. Seymour. Disjoint paths—a survey. SIAM Journal on Algebraic and Discrete Methods, 6(2):300– 305, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Rubinstein:1986:NRQ

[Rub86] Zalman Rubinstein. On the number of real quadratic factors of polynomials. SIAM Journal on Algebraic and Discrete Methods, 7(1):53–59, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Ruskey:1980:ASB

[Rus80] Frank Ruskey. On the average shape of binary trees. SIAM Journal on Algebraic and Discrete Methods, 1(1):43–50, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Redheffer:1983:MPC

[RV83] Ray Redheffer and Alexander Voigt. Matrix products that can be evaluated in closed form. SIAM Journal on Algebraic and Discrete Methods, 4(2):223–230, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Redheffer:1982:CMC

[RZ82] Ray Redheffer and Zhi Ming Zhou. A class of matrices connected with Volterra preypredator equations. SIAM Journal on Algebraic and Discrete Methods, 3(1):122–134, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Saks:1980:DNI

[Sak80] Michael Saks. Dilworth numbers, incidence maps and product partial orders. SIAM Journal on Algebraic and Discrete Methods, 1(2):211–215, June 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Saks:1982:CPG

[Sak82] Michael Saks. A class of perfect graphs associated with planar rectilinear regions. SIAM Journal on Algebraic and Discrete Methods, 3(3):330–342, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Sallee:1984:MCT

[Sal84] John F. Sallee. The middlecut triangulations of the n-cube. SIAM Journal on Algebraic and Discrete Methods, 5(3):407–419, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Sampson:1980:NCD

[Sam80] Allan R. Sampson. Nonnegative Cholesky decomposition and its application to association of random variables. SIAM Journal on Algebraic and Discrete Methods, 1(3):284–291, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Saxe:1980:DPA

[Sax80]  James B. Saxe.  Dynamic-

programming algorithms for recognizing small-bandwidth graphs in polynomial time. SIAM Journal on Algebraic and Discrete Methods, 1(4):363–369, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Shinoda:1983:CTC

[SCC83] Shoji Shinoda, Wai Kai Chen, and Shu Park Chan. The concept of two-chord tiesets and its application to an algebraic characterization of non-series-parallel graphs. SIAM Journal on Algebraic and Discrete Methods, 4(2):244–252, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Shao:1987:ESP

[Sha87] Jia Yu Shao. The exponent set of primitive, nearly reducible matrices. SIAM Journal on Algebraic and Discrete Methods, 8(4):578–584, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Shepp:1980:FIS

[She80] L. A. Shepp. The FKG inequality and some monotonicity properties of partial orders. SIAM Journal on Algebraic and Discrete Methods, 1(3):295– 299, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Shearer:1981:CPC

[She81] James B. Shearer. A counterexample to a bin packing conjecture. SIAM Journal on Algebraic and Discrete Methods, 2(3):309–310, September 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Shearer:1982:CPG

[She82] James B. Shearer. A class of perfect graphs. SIAM Journal on Algebraic and Discrete Methods, 3(3):281–284, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Shrager:1984:OSC

[Shr84] Richard I. Shrager. Optical spectra from chemical titration: an analysis by SVD. SIAM Journal on Algebraic and Discrete Methods, 5(3):351–358, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Satyanarayana:1986:IGR

[SK86] A. Satyanarayana and Zohel Khalil. On an invariant of graphs and the reliability polynomial. SIAM Journal on Algebraic and Discrete Methods, 7(3):399–403, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Seroussi:1983:SRF

[SL83] G. Seroussi and A. Lempel. On symmetric representations of finite fields. SIAM Journal on Algebraic and Discrete Methods, 4(1):14–21, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Smith:1981:MWI

[Smi81] Ronald L. Smith. M-matrices whose inverses are stochastic. SIAM Journal on Algebraic and Discrete Methods, 2(3):259– 265, September 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Sawami:1985:OOP

[SN85] Hideo Sawami and Hiroshi Niki. Optimum overrelaxation parameter for the SOR method for solving the eigenvalue problem. SIAM Journal on Algebraic and Discrete Methods, 6(2):278–282, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Simoes-Pereira:1984:NOS

[SP84] J. M. S. Simões-Pereira. A note on optimal and suboptimal digraph realizations of quasidistance matrices. SIAM Journal on Algebraic and Discrete Methods, 5(1):117–132, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Schmidt:1985:KGS

[SS85] Günther Schmidt and Thomas Ströhlein. On kernels of graphs and solutions of games: a synopsis based on relations and fixpoints. SIAM Journal on Algebraic and Discrete Methods, 6 (1):54–65, January 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Searle:1984:NHL

[SSM84] Shayle R. Searle, William H.

Swallow, and Charles E. McCulloch. Nontestable hypotheses in linear models. SIAM Journal on Algebraic and Discrete Methods, 5(4):486–496, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Soun:1980:SCR

[ST80] Y. Soun and K. Truemper.

Single commodity representation of multicommodity networks. SIAM Journal on Algebraic and Discrete Methods, 1(3):348–358, 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Stanley:1980:WGH

[Sta80a]  Richard P. Stanley.  Weyl

groups, the hard Lefschetz Theorem, and the Sperner property. SIAM Journal on Algebraic and Discrete Methods, 1(2):168–184, June 1980. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Stanton:1980:SEK

[Sta80b] Dennis Stanton. Some Erdős– Ko–Rado theorems for Chevalley groups. SIAM Journal on Algebraic and Discrete Methods, 1(2):160–163, June 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Stahl:1985:HTI

[Sta85] Saul Stahl. Homology as a tool in integer programming. SIAM Journal on Algebraic and Discrete Methods, 6(4):641–642, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Stensholt:1986:KCF

[Ste86] Eivind Stensholt. Keldysh chains and factorizations of matrix polynomials. SIAM Journal on Algebraic and Discrete Methods, 7(3):462–469, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Sugihara:1983:SPD

[Sug83] Kōkichi Sugihara. On some problems in the design of plane skeletal structures. SIAM Journal on Algebraic and Discrete Methods, 4(3):355–362, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Simion:1986:DPO

[SW86] Rodica Simion and Herbert S. Wilf. The distribution of prefix overlap in consecutive dictionary entries. SIAM Journal on Algebraic and Discrete Methods, 7(3):470–475, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Shier:1987:AMA

[SW87] Douglas R. Shier and David E.

Whited. Algebraic methods applied to network reliability problems. SIAM Journal on Algebraic and Discrete Methods, 8(2):251–262, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Shapiro:1983:RSM

[SWG83] Louis W. Shapiro, Wen Jin Woan, and Seyoum Getu. Runs, slides and moments. SIAM Journal on Algebraic and Discrete Methods, 4(4):459–466, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Szpankowski:1987:SLR

[Szp87] Wojciech Szpankowski. Solution of a linear recurrence equation arising in the analysis of some algorithms. SIAM Journal on Algebraic and Discrete Methods, 8(2):233–250, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Takacs:1981:RFR

[Tak81] Lajos Takács. Random flights on regular polytopes. SIAM Journal on Algebraic and Discrete Methods, 2(2):153–171, June 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Tamir:1983:CBM

[Tam83] Arie Tamir. A class of balanced matrices arising from location problems. SIAM Journal on Algebraic and Discrete Methods, 4(3):363–370, September 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Tannenbaum:1983:P

[Tan83] Peter Tannenbaum. Partitions of Z_{2}^{n}. SIAM Journal on Algebraic and Discrete Methods, 4(1):22– 29, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Tanner:1984:ECG

[Tan84] R. Michael Tanner. Explicit concentrators from generalized N-gons. SIAM Journal on Algebraic and Discrete Methods, 5(3):287–293, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Tarjan:1985:ACC

[Tar85] Robert Endre Tarjan. Amortized computational complexity. SIAM Journal on Algebraic and Discrete Methods, 6(2):306– 318, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Taylor:1982:SCC

[Tay82] R. Taylor. Switchings constrained to 2-connectivity in simple graphs. SIAM Journal on Algebraic and Discrete Methods, 3 (1):114–121, March 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Tazawa:1985:DCM

[Taz85] Shinsei Tazawa. Decomposition of a complete multipartite graph into isomorphic claws. SIAM Journal on Algebraic and Discrete Methods, 6(3):413–417, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Tchuente:1985:PFS

[Tch85] Maurice Tchuente. Permutation factorization on star-connected networks of binary automata. SIAM Journal on Algebraic and Discrete Methods, 6(3):537–540, July 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Tiersma:1985:SGS

[Tie85] H. J. Tiersma. On subcodes of generalized second order Reed– Muller codes. SIAM Journal on Algebraic and Discrete Methods, 6(4):583–591, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Toczylowski:1984:HRI

[Toc84] Eugeniusz Toczyłowski. A hierarchical representation of the inverse for sparse matrices. SIAM Journal on Algebraic and Discrete Methods, 5(1):43–56, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Todd:1984:COM

[Tod84a]  Michael J. Todd.  Comple-

mentarity in oriented matroids. SIAM Journal on Algebraic and Discrete Methods, 5(4):467–485, 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Todd:1984:NT

[Tod84b] Michael J. Todd. J′: a new triangulation of R^{n}. SIAM Journal on Algebraic and Discrete Methods, 5(2):244–254, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Togai:1986:ASV

[Tog86] Masaki Togai. An application of the singular value decomposition to manipulability and sensitivity of industrial robots. SIAM Journal on Algebraic and Discrete Methods, 7(2):315–320, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Tovey:1985:HCM

[Tov85] Craig A. Tovey. Hill climbing with multiple local optima. SIAM Journal on Algebraic and Discrete Methods, 6(3):384–393, July 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Trench:1985:EIF

[Tre85] William F. Trench. Explicit inversion formulas for Toeplitz band matrices. SIAM Journal on Algebraic and Discrete Methods, 6(4):546–554, 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Trench:1986:EPC

[Tre86] William F. Trench. On the eigenvalue problem for a class of band matrices including those with Toeplitz inverses. SIAM Journal on Algebraic and Discrete Methods, 7(2):167–179, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Tran:1985:DRM

[TS85] Minh Thanh Tran and Mahmoud E. Sawan. On the discrete Riccati matrix equation. SIAM Journal on Algebraic and Discrete Methods, 6 (1):107–108, January 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Tanny:1983:STC

[TZ83] Stephen M. Tanny and Michael Zuker. Signs of terms in a combinatorial recursion. SIAM Journal on Algebraic and Discrete Methods, 4(4):437–441, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Uhlig:1985:FVC

[Uhl85] Frank Uhlig. The field of values of a complex matrix, an explicit description in the 2 × 2 case. SIAM Journal on Algebraic and Discrete Methods, 6(4):541– 545, 1985. CODEN SJAMDU.

ISSN 0196-5212 (print), 2168-345X (electronic).

Ursic:1983:ESS

[UP83] Silvio Ursic and Cyro Patarra. Exact solution of systems of linear equations with iterative methods. SIAM Journal on Algebraic and Discrete Methods, 4(1):111–115, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Varga:1982:FMC

[VC82] R. S. Varga and D. Y. Cai. On the LU factorization of Mmatrices: cardinality of the set P_{n}^{g}(A). SIAM Journal on Algebraic and Discrete Methods, 3(2):250–259, 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Vichniac:1984:IDA

[Vic84] Gérard Y. Vichniac. Instability in discrete algorithms and exact reversibility. SIAM Journal on Algebraic and Discrete Methods, 5(4):596–602, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Victory:1985:NSM

[Vic85] H. D. Victory, Jr. On nonnegative solutions of matrix equations. SIAM Journal on Algebraic and Discrete Methods, 6(3):406–412, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Vo:1981:SCL

[Vo81] Kiem Phong Vo. The Schensted correspondence and lexicographic matchings on multisubset lattices. SIAM Journal on Algebraic and Discrete Methods, 2(3):324–332, September 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

vonKaenel:1981:SCC

[von81] Pierre A. von Kaenel. Some constructions for convolutional and block codes. SIAM Journal on Algebraic and Discrete Methods, 2(2):121–126, June 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Walsh:1983:GNM

[Wal83] T. R. Walsh. Generating nonisomorphic maps without storing them. SIAM Journal on Algebraic and Discrete Methods, 4(2):161–178, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Walter:1985:CSC

[Wal85] Gilbert G. Walter. Complexity and stability in compartmental models. SIAM Journal on Algebraic and Discrete Methods, 6 (1):39–46, January 1985. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Walter:1986:AM

[Wal86] Colin D. Walter. Adjacency matrices. SIAM Journal on Algebraic and Discrete Methods, 7(1):18–29, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Wang:1983:PG

[Wan83] Kai Wang. On the polynomials of graphs. SIAM Journal on Algebraic and Discrete Methods, 4(4):522–528, 1983. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Wang:1984:SSG

[Wan84] Kai Wang. Spectra of some graphs. SIAM Journal on Algebraic and Discrete Methods, 5(1):57–60, 1984. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Waterhouse:1987:SMC

[Wat87] William C. Waterhouse. The structure of monomial circulant matrices. SIAM Journal on Algebraic and Discrete Methods, 8(3):467–482, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Wei:1986:BLP

[WCLO86] W. D. Wei, Y. Z. Cai, C. L. Liu, and A. M. Odlyzko. Balloting labelling and personnel assignment. SIAM Journal on Algebraic and Discrete Methods, 7(1):150–158, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Wendel:1987:GMF

[Wen87] H. Wendel. The general minimum fill-in problem. SIAM Journal on Algebraic and Discrete Methods, 8(4):710–745, 1987. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

West:1982:GDT

[Wes82]  Douglas B. West.  Gossip-

ing without duplicate transmissions. SIAM Journal on Algebraic and Discrete Methods, 3(4):418–419, December 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

White:1986:PAN

[Whi86] R. E. White. Parallel algorithms for nonlinear problems. SIAM Journal on Algebraic and Discrete Methods, 7(1):137–149, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Wilf:1986:NMI

[Wil86] Herbert S. Wilf. The number of maximal independent sets in a tree. SIAM Journal on Algebraic and Discrete Methods, 7(1):125– 130, 1986. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Wimmer:1982:ARE

[Wim82] H. K. Wimmer. The algebraic Riccati equation without complete controllability. SIAM Journal on Algebraic and Discrete Methods, 3(1):1–12, March 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Wimmer:1984:DBI

[Wim84] Harald K. Wimmer. On the dimension of bi-infinite systems. SIAM Journal on Algebraic and Discrete Methods, 5(1):9–10, March 1984. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Winkler:1983:CAP

[Win83] P. M. Winkler. Correlation among partial orders. SIAM Journal on Algebraic and Discrete Methods, 4(1):1–7, March 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Wong:1980:WCA

[Won80] Richard T. Wong. Worst-case analysis of network design problem heuristics. SIAM Journal on Algebraic and Discrete Methods, 1(1):51–63, 1980. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Wisniewski:1982:PAN

[WS82] John A. Wisniewski and Ahmed H. Sameh. Parallel algorithms for network routing problems and recurrences. SIAM Journal on Algebraic and Discrete Methods, 3(3):379–394, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

West:1981:SUC

[WT81] Douglas B. West and Craig A. Tovey. Semiantichains and unichain coverings in direct products of partial orders. SIAM Journal on Algebraic and Discrete Methods, 2(3):295–305, September 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

White:1983:AGS

[WW83] Neil L. White and Walter Whiteley. The algebraic geometry of stresses in frameworks. SIAM Journal on Algebraic and Discrete Methods, 4(4):481–511, 1983. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

White:1987:AGM

[WW87] Neil White and Walter Whiteley. The algebraic geometry of motions of bar-and-body frameworks. SIAM Journal on Algebraic and Discrete Methods, 8(1):1–32, 1987. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Yannakakis:1981:CMF

[Yan81] Mihalis Yannakakis. Computing the minimum fill-in is NPcomplete. SIAM Journal on Algebraic and Discrete Methods, 2 (1):77–79, March 1981. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Yannakakis:1982:CPO

[Yan82] Mihalis Yannakakis. The complexity of the partial order dimension problem. SIAM Journal on Algebraic and Discrete Methods, 3(3):351–358, 1982. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Yao:1982:SDP

[Yao82] F. Frances Yao. Speed-up in dynamic programming. SIAM Journal on Algebraic and Discrete Methods, 3(4):532–540, December 1982. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Zemanian:1981:ESP

[Zem81] A. H. Zemanian. Equilibrium and stability in a periodic marketing ring. SIAM Journal on Algebraic and Discrete Methods, 2(3):275–288, September 1981. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Zemel:1985:PAG

[Zem85] Eitan Zemel. Probabilistic analysis of geometric location problems. SIAM Journal on Algebraic and Discrete Methods, 6(2):189–200, 1985. CODEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).